### v1

# Position: LLMs Can be Good Tutors in Foreign Language Education

Abstract
While recent efforts have begun integrating large language models (LLMs) into foreign language education (FLE), they often rely on traditional approaches to learning tasks without fully embracing educational methodologies, thus lacking adaptability to language learning. To address this gap, we argue that LLMs have the potential to serve as effective tutors in FLE. Specifically, LLMs can play three critical roles: (1) as data enhancers, improving the creation of learning materials or serving as student simulations; (2) as task predictors, serving as learner assessment or optimizing learning pathway; and (3) as agents, enabling personalized and inclusive education. We encourage interdisciplinary research to explore these roles, fostering innovation while addressing challenges and risks, ultimately advancing FLE through the thoughtful integration of LLMs.
Machine Learning, ICML
1 Introduction
Foreign Language Education (FLE) has long been a cornerstone of global education and a critical component of K-12 curricula, equipping students with the linguistic and cultural competencies necessary for an interconnected world. However, traditional FLE methods often fall short in addressing the diverse needs of learners (Hou, 2020). Challenges such as limited personalization, scalability constraints, and the lack of real-time feedback are particularly pronounced in large classroom settings (Ehrenberg et al., 2001), where individual attention is scarce, leaving many students disengaged or struggling to keep pace. Addressing these shortcomings requires innovative approaches that not only enhance the quality of instruction but also adapt to the unique learning trajectories of students (Eaton, 2010).
In parallel, the advent of LLMs has opened new possibilities for transforming FLE (Caines et al., 2023). These models, powered by large-scale pre-training, exhibit remarkable capabilities in natural language understanding and generation, making them promising candidates for roles traditionally filled by human tutors. By leveraging LLMs as AI tutors, we may overcome many of the inherent limitations of conventional teaching methods, offering scalable, interactive, and personalized learning experiences (Chen et al., 2024; Schmucker et al., 2024). Therefore, this position paper argues that LLMs can be effective tutors in foreign language education, complementing human expertise and addressing key limitations of traditional methods.
Figure 1: Involved disciplines of LLM for FLE.
Figure 2: Overview of roles of LLMs in foreign language education. An overview of related literature is provided in Appendix A.
As shown in Figure 1, FLE intersects with multiple disciplines, each of which underscores the potential of LLMs to revolutionize this domain. From the perspective of (1) computer science, advancements in machine learning and NLP have enabled LLMs to process and generate human-like language at an unprecedented scale; (2) linguistics (Radford et al., 2009) contributes a deeper understanding of grammar, phonetics, and semantics, allowing LLMs to generate accurate and understandable language outputs; (3) education provides the foundation for designing effective pedagogical strategies, ensuring that LLMs can deliver personalized, engaging, and developmentally appropriate learning experiences; and finally, (4) psycholinguistics (Steinberg & Sciarini, 2013) bridges the gap between language acquisition and cognitive processes, enabling LLMs to optimize learner interactions by adapting to individual needs and fostering meaningful engagement. Together, these disciplines position LLMs as uniquely capable of addressing the multifaceted challenges of foreign language education.
Moreover, FLE encompasses four core skills: listening, speaking, reading, and writing, each of which can be significantly enhanced by LLMs. For listening, LLMs can generate diverse audio materials (Ghosal et al., 2023) and facilitate interactive voice-based exercises, helping learners improve their ability to discern pronunciation, intonation, and contextual meaning. In speaking, LLMs can simulate realistic conversations (Siyan et al., 2024), provide pronunciation feedback, and scaffold learners’ oral communication skills through iterative practice. For reading, LLMs can curate leveled texts, generate comprehension questions (Samuel et al., 2024), and engage learners in discussions that deepen their understanding of written content. Finally, in writing, LLMs can offer real-time grammar, syntax, and style feedback while assisting with idea generation and iterative revisions (Stahl et al., 2024a). By addressing these core skills holistically, LLMs have the potential to deliver a comprehensive and adaptive learning experience.
Despite these opportunities, the deployment of LLMs in FLE must be approached carefully, ensuring that their integration complements rather than replaces human tutors (Jeon & Lee, 2023). As illustrated in Figure 2, this paper explores three critical roles of LLMs in this context: their function as data enhancers (Section 4) to optimize learning materials, their capacity as task predictors (Section 5) to tailor educational solutions, and their potential as agents (Section 6) that deliver interactive and adaptive language instruction. By examining these roles, we aim to demonstrate how LLMs can address the limitations of traditional FLE methods while advancing our understanding of intelligent tutoring systems. Additionally, we discuss potential challenges, ethical considerations, and future directions (Section 7) for integrating LLMs into FLE, offering a technical guideline for researchers and educators to harness their transformative potential. We also describe the paradigm shift of leveraging AI for FLE, starting from the last century, as one of our contributions in Section 3.
Figure 3: Roadmap of Foreign Language Education.
2 Background
2.1 Foreign Language Education
FLE (Watzke, 2003; Rashov, 2024) has long been a cornerstone of global communication and cultural exchange. Traditional methods often emphasize grammar rules, vocabulary memorization, and repetitive practice, supplemented by limited opportunities for real-world application. Such approaches are often constrained by the availability of skilled teachers, the diversity of learners’ needs, and the lack of personalized feedback (Williams et al., 2004). To bridge the gaps, many technologies for FLE have been proposed (Alhusaiyan, 2024), focusing on solving specific tasks instead of describing the whole picture of foreign language tutoring. While intelligent language tutoring systems (Schwind, 1990; C Angelides & Garcia, 1993) have the potential to create adaptive environments, attention to this field is relatively less compared to other subjects like science and mathematics. One key reason lies in the inherent complexity of language as an ill-defined domain (Schmidt & Strasser, 2022), posing a great challenge in establishing a valid automatic analysis of learner languages due to the vast variability and unpredictability of human language.
2.2 Large Language Models
The potential of LLMs in education (Alhafni et al., 2024), particularly in FLE (Gao et al., 2024; Karataş et al., 2024), is immense. Benefiting from large-scale pre-training on extensive corpora, LLMs have demonstrated emergent abilities including (1) in-context learning (Dong et al., 2022), which allows the model to adapt to new tasks and provide contextually relevant responses based on a few examples provided during the interaction; (2) instruction following (Zeng et al., 2024), which enables the model to process and execute complex user instructions with high accuracy; and (3) reasoning and planning (Huang et al., 2024b), which allows the model to generate coherent, structured, and context-aware outputs, even for tasks that require multi-step thinking.
However, these fundamental capabilities, while impressive, are not sufficient to fully meet the unique demands of FLE. Teaching a foreign language requires more than generating grammatically correct sentences or providing accurate translations; it demands a nuanced understanding of pedagogy, learner psychology, and cultural context. Maurya et al. (2024) propose an evaluation taxonomy that identifies eight critical dimensions for assessing AI tutors. These dimensions can be broadly categorized into two groups. (1) Problem-solving abilities assess the technical capabilities of LLMs to perform tasks relevant to FLE. (2) Pedagogical alignment abilities evaluate how well the LLM aligns with effective teaching and learning principles. Pedagogical alignment includes the model’s ability to adapt to the learner’s proficiency level, provide scaffolded feedback, foster engagement, and maintain motivation. While LLMs can give direct answers, their ability to replicate these nuanced teaching strategies remains a challenge (Wang et al., 2024a). Additionally, pedagogical alignment requires sensitivity to cultural and contextual factors, ensuring that the learning experience is relevant and meaningful to the student.
3 Paradigm Shift
The development of AI models for FLE can be broadly categorized into four successive generations as shown in Figure 3: (1) rule-based models, (2) statistical models, (3) neural models, and (4) large language models. We leave the detailed description in Appendix B.
Our position. We foresee next-generation LLMs with deeper alignment to pedagogical principles and stronger guardrails to mitigate misinformation and bias. Future models may integrate multimodal data (e.g., text, image, video, speech) and employ meta-learning to adapt in real time to diverse learner profiles. These improvements will reinforce the position that LLMs can evolve into even more effective tutors for FLE.
4 LLMs as Data Enhancers
Education is a high-stake area where any form of LLMs’ hallucination could cause devastating harm to humans’ cognition activities (Ho et al., 2024). One of the hallucination causes is from data (Huang et al., 2023). Therefore, high-quality and diverse data resources (Long et al., 2024) are critical to ensuring the reliability of incorporating LLMs into FLE. The 1) creation, 2) reformation, and 3) annotation of educational materials are crucial to delivering effective and engaging teaching. Traditional resource development methods often lack the scalability, adaptability, and personalization necessary to meet the diverse needs of learners (Feng et al., 2021; Shorten et al., 2021). In contrast, LLMs emerge as transformative tools capable of enhancing these processes (Wang et al., 2024c; Liu et al., 2024c). This section explores how LLMs serve as data enhancers in FLE.
4.1 Data Creation
Creating pedagogically sound and learner-specific data is a cornerstone of personalized learning. However, manually creating such resources is time-consuming and often fails to address the wide range of learner needs (Cochran et al., 2022). LLMs can revolutionize this process by generating tailored and diverse educational content or responses on demand (Zha et al., 2023; Cochran et al., 2023).
Educational Materials Generation.
A primary use of LLMs in data creation is the generation of educational questions aligned with specific learning objectives. Due to their superior contextual understanding, classic rule-based approaches have largely been eclipsed by neural network-based techniques (Kurdi et al., 2020; Rathod et al., 2022; Mulla & Gharpure, 2023). LLMs can produce answer-aware (whose target answer is known) or answer-agnostic (whose answer is open) (Zhang et al., 2021), resulting in more nuanced exercises and assessments (Xiao et al., 2023).
Student Simulation.
Simulating the learner’s perspective is crucial for designing adaptive instructional materials. Traditional surveys and standardized tests often fail to capture the complexity of dynamic learner behaviors (Käser & Alexandron, 2024). In contrast, LLM-based approaches enable high-fidelity, context-aware student simulations (Liu et al., 2024d; Yue et al., 2024), generating synthetic learners who exhibit realistic mastery levels and evolving behaviors. For instance, Generative Students (Lu & Wang, 2024) create simulated learners with various competency levels, while EduAgent (Xu et al., 2024) integrates cognitive priors to model complex learning trajectories and behaviors better.
Discussion.
While LLMs excel at generating educational content, current approaches mainly focus on question creation, leaving many areas of FLE underexplored. Essential tasks like generating culturally rich reading materials, context-dependent writing prompts, or dynamic comprehension exercises are still lacking in diversity and depth. Additionally, the student simulations created by LLMs often fail to reflect long-term learning trajectories or the intricacies of individual learning progress, leading to challenges in creating truly adaptive systems.
4.2 Data Reformation
In addition to creating new content, LLMs can adapt existing materials to better align with current needs. This process, commonly referred to as data reformation, involves (1) changing data types or modalities, (2) paraphrasing materials to match learner proficiency, and (3) enriching raw data with auxiliary signals or contextual content.
Teaching Material Transformation.
Transforming existing materials into different forms can yield more comprehensive and immersive learning experiences. For example, Book2Dial(Wang et al., 2024b) generates teacher-student dialogues grounded in textbooks, keeping the content both relevant and informative. Their approach includes multi-turn question generation and answering(Kim et al., 2022), dialogue inpainting (Dai et al., 2022), and role-playing. Likewise, Slide2Lecture (Zhang-Li et al., 2024) automatically converts lecture slides into structured teaching agendas, enabling interactive follow-up and deeper learner engagement.
Simplification and Paraphrasing.
Another vital application is simplifying or paraphrasing complex texts to specified readability levels (Huang et al., 2024a) without losing key concepts (Al-Thanyyan & Azmi, 2021). This is particularly beneficial in FLE settings, where language beginners often face advanced vocabulary and complex structures (Day et al., 2025). Recent advancements in controllable generation (Zhang et al., 2023) leverage model fine-tuning on curated datasets (Zeng et al., 2023) or decoding-time interventions (Liang et al., 2024), thereby allowing educators to specify text complexity, style, or tone.
Cultural Context Adaptation.
Beyond linguistic correctness, cultural nuance is pivotal in FLE (Byram, 1989, 2008). LLMs can facilitate this process by recontextualizing existing materials to reflect the cultural and social norms of the target language (Liu et al., 2024a). For instance, a short story originally set in an English-speaking environment may be adapted for French or Japanese learners by adjusting the characters’ names, idiomatic expressions, or social customs, while still preserving core instructional goals. This cultural adaptation not only enhances learner engagement but also strengthens cross-cultural competencies.
Discussion.
While LLM-based data reformation can significantly enhance FLE, several gaps warrant attention. Most current studies prioritize textual forms or single-modal approaches, which may overlook valuable multimodal resources such as interactive video and audio-based content (Ghosal et al., 2023). Furthermore, cultural adaptation, although promising, remains underexplored in practical classroom scenarios, particularly for underrepresented personas and culturally sensitive topics. AlKhamissi et al. (2024) demonstrate how cultural misalignment can increase bias, suggesting Anthropological Prompting to strengthen cultural alignment. However, robust empirical evaluations are still limited across diverse learners and linguistic backgrounds. These issues highlight the importance of interdisciplinary research, where educators, linguists, and technologists collaborate to refine data reformation techniques that broaden inclusivity and effectiveness in FLE.
4.3 Data Annotation
While Data Creation focuses on generating learner-specific data, it often prioritizes diversity and adaptability over precision. The approach is particularly useful for tasks with large label spaces (Ding et al., 2024). In contrast, Data Annotation emphasizes producing high-quality, meticulously labeled data that is essential for tasks requiring accuracy and consistency. Unlike data creation, annotated data often undergoes rigorous validation to ensure its accuracy and relevancy (Artemova et al., 2024).
Annotation Generation.
LLMs can be central to generating a variety of annotations, including categorical labels, rationales, pedagogical feedback, and linguistic features such as discourse relations. Recent prompt engineering and fine-tuning techniques have further expanded LLMs’ annotation capabilities. For instance, Ye et al. (2024) leverage GPT-4 to annotate structured explanations for Chinese grammatical error correction, while Samuel et al. (2024) examine GPT-4 as a surrogate for human annotators in low-resource reading comprehension tasks. Likewise, Siyan et al. (2024) deploy GPT-4-Turbo for audio transcript annotations. However, inconsistencies across LLMs (Törnberg, 2024) remain a serious challenge, posing risks to educational reliability.
Annotation Assessment.
Although LLM-based annotation generation is efficient, it also raises critical issues of bias, calibration, and validity—particularly in low-resource language contexts (Bhat & Varma, 2023; Jadhav et al., 2024). Automated or semi-automated evaluation strategies have emerged to address these quality concerns. For example, LLMs-as-Judges (Li et al., 2024a, b; Gu et al., 2024) reduce human overhead by automating evaluation, an approach increasingly explored in education-focused applications (Chiang et al., 2024; Zhou et al., 2024). However, purely automated frameworks can still propagate errors or bias.
Discussion.
Although LLMs provide efficient data annotation, the inconsistency across different models remains a critical concern, affecting the quality and reliability of annotated educational materials. These discrepancies hinder the creation of universally reliable FLE content, especially in diverse linguistic and cultural contexts. Additionally, automated annotations often lack the nuance needed for pedagogical applications, making it essential to involve human oversight in critical cases to mitigate errors or biases.
Our position. We acknowledge the current limitations in LLM-based data creation, reformation, and annotation for FLE. However, we believe that with continued interdisciplinary collaboration, including input from educators, linguists, and technologists, these challenges can be addressed. Future advancements should focus on enhancing the accuracy and diversity of generated content, improving multimodal and culturally sensitive learning materials, and integrating more robust systems for human-LLM collaboration (Li et al., 2023; Wang et al., 2024e) in data annotation. This approach will ensure that LLMs can fully realize their potential as effective tutors in FLE.
5 LLMs as Task Predictors
Task-Based Language Learning (TBLL) (Nunan, 1989; Willis, 2021) as a methodological approach is one of the critical factors of modern foreign language teaching. LLMs have demonstrated remarkable capabilities in understanding and generating human language, making them well-suited for addressing numerous tasks in FLE. These tasks can be broadly categorized into three types based on their nature and the role of LLMs: 1) Discriminative, 2) Generative, and 3) Mixed of the above two roles.
5.1 Discriminative Task Predictors
Discriminative tasks in FLE primarily involve classifying learner inputs or grading their future performance. Below are some applications that are still calling for improvements:
Automated Assessment.
The task aims to automatically grade students’ assignments, including essay scoring (Seßler et al., 2024; Li & Liu, 2024; Syamkumar et al., 2024), short answer grading (Schneider et al., 2023; Henkel et al., 2024), and spoken language evaluation (Gao et al., 2023; Fu et al., 2024). LLMs can process learners’ submissions to judge grammar, lexical diversity, coherence, and even spoken fluency, providing instant feedback. This scalability is particularly appealing for large classes, where human evaluators are often overwhelmed and unable to provide timely, personalized critique (Mizumoto & Eguchi, 2023).
Knowledge Tracing.
Given sequences of learning interactions in online learning systems, Knowledge Tracing identifies and tracks students’ evolving mastery of target skills (Shen et al., 2024b; Xu et al., 2023). LLM-based methods have been explored in cold-start scenarios (Zhan et al., 2024; Jung et al., 2024), offering strong generalization by inferring latent learner states from limited data. These approaches can support adaptive learning pathways, giving personalized recommendations based on predicted performance and knowledge gaps.
Discussion.
Despite their promise in automating and personalizing these discriminative tasks, LLMs still grapple with notable limitations that hinder their utility as robust tutoring tools. First, misalignment of assessment with expert instructors poses risks: machine-generated scores may deviate from established rubrics or neglect qualitative nuances, leading to potential discrepancies in grading quality (Kundu & Barbosa, 2024). Second, the lack of empathy compounds this issue, as assessments devoid of human judgment risk discouraging learners or overlooking subtle motivational factors (Sharma et al., 2024). Knowledge tracing approaches, while promising in cold-start scenarios, struggle with capturing the complexity of long-term learning trajectories and deeper cognitive processes (Cho et al., 2024). These concerns point to the need for more transparent and human-centered methods in utilizing LLMs for assessment.
5.2 Generative Task Predictors
Generative tasks involve producing new content or responses. LLMs are known to be adept in these tasks due to their natural language generation capabilities.
Grammatical Error Correction and Explanation.
In foreign language writing, errors often reveal learners’ gaps in grammar and vocabulary (Hyland & Hyland, 2006). LLMs can detect and correct these errors (Bryant et al., 2023; Ye et al., 2023), offering concise explanations (Ye et al., 2024) that reinforce language rules. By streamlining error detection and pedagogically framing corrections, learners deepen their linguistic understanding.
Feedback Generation.
Quizzes and exercises remain vital in FLE for practice and targeted remediation (Rashov, 2024). LLMs enhance this process by delivering prompt, personalized feedback that pinpoints strengths and addresses weaknesses (Borges et al., 2024). This scalability enables learners to self-regulate and refine their skills without relying solely on human graders (Stamper et al., 2024).
Socratic Dialogue.
Moving beyond straightforward Q&A, Socratic questioning promotes critical thinking and self-reflection (Paul & Elder, 2007). SocraticLM (Liu et al., 2024b), for example, aligns an LLM with open-ended, inquiry-based teaching principles, guiding learners through iterative exploration rather than prescriptive correction. In theory, this fosters deeper conceptual understanding and active learner engagement.
Discussion.
Despite the promise of LLM-based generation in FLE, multiple uncertainties persist. Determining how to provide automatic feedback that genuinely maximizes learning outcomes is an ongoing challenge (Stamper et al., 2024), particularly given education’s risk-averse culture and high accountability standards (Xiao et al., 2024). Moreover, while LLMs like SocraticLM have demonstrated success in domains like mathematics, their applicability to FLE contexts has not been thoroughly validated (Liu et al., 2024b). As such, the design of strategies and follow-up queries remains an open question in ensuring that these systems track and respond to learners’ cognitive states effectively.
5.3 Mixed Task Predictors
Mixed tasks integrate discriminative and generative elements, requiring LLMs to evaluate learner inputs and generate meaningful feedback or suggestions. These tasks are particularly valuable in fostering an interactive and adaptive learning experience, as they bridge the gap between evaluation and instruction.
Automated Assessment with Feedback.
While discriminative systems for essay scoring and speech evaluation primarily focus on assigning grades, LLMs extend these capabilities by simultaneously generating formative feedback (Katuka et al., 2024; Stahl et al., 2024b). For example, an LLM can evaluate the coherence and lexical diversity of a written assignment, then offer specific revision strategies. In speaking practice, it can measure fluency and pronunciation accuracy while suggesting drills to refine intonation or stress patterns. Through this combination of scoring and tailored advice, learners gain a deeper understanding of their strengths and areas for improvement.
Error Analysis.
Error Analysis systematically uncovers and categorizes learners’ missteps, from syntactic lapses in writing to flawed pronunciations in speaking (James, 2013; Erdoğan, 2005). LLMs functioning in a mixed capacity can classify these errors and generate corrective guidance, providing revised sentences, clarifications of grammatical rules, or remediation exercises for identified weaknesses (Myles, 2002; Mashoor & Abdullah, 2020). Such insight facilitates targeted interventions that enhance language proficiency across modalities, including reading and listening.
Discussion.
Mixed-task systems hold promise by combining assessment and feedback generation, but they face notable challenges. One major issue is the weak alignment between scoring mechanisms and the quality of feedback provided (Stahl et al., 2024b). For example, while essay scoring systems may deliver comprehensive evaluations, the feedback often lacks specificity, limiting its instructional value. Additionally, although error analysis has potential, the absence of standardized pedagogical benchmarks, especially in oral tasks, hampers the reliability and comparability of LLM-based tools (Leu Jr, 1982). Furthermore, by addressing errors across these modalities, LLMs can deliver a holistic and adaptive learning experience that fosters deeper language proficiency (Zhao et al., 2024).
Our position. While LLMs offer scalable solutions for task prediction in FLE, their current limitations—such as misalignment with expert assessments, lack of empathy, and weak alignment between assessment and feedback—require ongoing refinement. Future research should focus on improving model transparency, enhancing the cultural and emotional sensitivity of LLMs, and refining task predictors to better reflect long-term learning trajectories and learner motivation. Additionally, developing standardized pedagogical benchmarks for error analysis, especially in oral tasks, will help ensure the consistency and reliability of LLM-generated feedback.
6 LLM-empowered Agent
In this section, we delve into the potential of LLMs as intelligent agents in FLE. LLMs can act as catalysts for personalized learning, addressing the long-standing scalability, adaptability, and inclusivity challenges in traditional teaching paradigms.
6.1 Fundamental Abilities
This section highlights five key abilities of LLM-empowered agents that enable them to function as adaptive tutors.
Knowledge Integration.
LLMs excel at merging structured educational knowledge graphs (Abu-Rasheed et al., 2024; Hu & Wang, 2024) with unstructured textual data (Li et al., 2024c; Modran et al., 2024), providing rich, contextualized information on linguistic constructs and cultural nuances. Their ability to perform real-time knowledge editing (Wang et al., 2024d; Zhang et al., 2024a) ensures learners receive content aligned with evolving language usage, addressing the inherent limitations of static materials.
Pedagogical Alignment.
LLMs require embedding with pedagogical principles to facilitate genuine learning experiences (Carroll, 1965; Taneja, 1995). Recent work incorporates theoretical frameworks, such as Bloom’s taxonomy (Bloom et al., 1956), to guide LLMs in systematically addressing different cognitive levels (Jiang et al., 2024b). Approaches like Pedagogical Chain of Thought (Jiang et al., 2024b) and preference learning (Sonkar et al., 2024; Rafailov et al., 2024) focus on aligning model responses with educational objectives.
Planning.
By assisting in crafting teaching objectives and lesson designs, LLMs can handle complex tasks such as differentiated instruction (Hu et al., 2024). LessonPlanner (Fan et al., 2024) has been proposed to assist novice teachers in preparing lesson plans, with expert interviews confirming its effectiveness. Zheng et al. (2024) propose a three-stage process to produce customized lesson plans, using Retrieval-Augmented Generation (RAG), self-critique, and subsequent refinement.
Memory.
Effective tutoring systems track learner histories and tailor subsequent interactions accordingly (Jiang et al., 2024a; Chen et al., 2024). When serving as memory-augmented agents, LLMs can retain individualized data—such as repeated grammar mistakes or overlooked vocabulary—thereby improving continuity and enabling consistent scaffolding of future learning tasks.
Tool Using.
Beyond textual interactions, LLM-based agents can integrate specialized tools to streamline the educational ecosystem, from cognitive diagnosis modules (Ma & Guo, 2019) to report generators (Zhou et al., 2025). By orchestrating these resources, LLMs seamlessly unify diverse utilities under a single interface, enhancing learner experience and instructional efficiency.
Discussion.
Existing research often overlooks the interplay among listening, speaking, reading, and writing in real-world language learning (Caines et al., 2023; Shetye, 2024). Most systems concentrate on text-oriented features, lacking robust benchmarks and methodologies to evaluate integrated, multimodal interactions. Recent work on pedagogical alignment (Razafinirina et al., 2024) largely addresses textual data, leaving out real-time speaking and listening dynamics that demand complex, rapid feedback. Similarly, while memory modules can track repeated written errors, their effectiveness in monitoring and improving learning efficacy remains underexplored.
6.2 Applications
Although still in its early stages, LLM-empowered agents have already started to show promising applications in FLE.
Classroom Simulation.
Classroom simulation leverages LLM-empowered agents to recreate complex, interactive learning settings without the logistical hurdles of organizing physical classrooms (Zhang et al., 2024b). By simulating virtual students and tutors, researchers can study pedagogical strategies at scale, generate diverse learner interactions, and refine teaching techniques. Moreover, this virtual data can be used to fine-tune LLMs for specific educational contexts and learner profiles (Liu et al., 2024b), offering a cost-effective and adaptable approach to language instruction.
Intelligent Tutoring System (ITS).
LLM-based agents have demonstrated the capacity to provide dynamic, personalized tutoring experiences (Kwon et al., 2024), effectively identifying learner weaknesses through large-scale linguistic analysis (Caines et al., 2023). This makes them promising for delivering individualized instruction at scale. Although current ITS applications in mathematics (Pal Chowdhury et al., 2024) and science (Stamper et al., 2024) have shown success, the extension to FLE requires nuanced handling of cultural and contextual elements, as well as the unpredictability of human language usage.
Discussion.
Despite the promise of these applications, critical challenges remain. Existing classroom simulation frameworks often lack standardized benchmarks for FLE, making it difficult to assess the efficacy and generalizability of developed systems (Zhang et al., 2024b). In addition, evaluating language-specific tutoring strategies, including real-time conversational practice and holistic skill integration, remains an underexplored frontier. Addressing these gaps requires new datasets and metrics centered on holistic skill development, as well as interdisciplinary collaboration.
Our position. We argue that to overcome current limitations, future research should focus on the integration of multimodal learning tasks (Sonlu et al., 2024) and the development of standardized frameworks for evaluating FLE simulations. Moreover, LLMs should evolve beyond text-based capabilities to provide real-time, context-sensitive feedback, particularly in speaking and listening. Interdisciplinary collaboration and the creation of new datasets tailored to FLE are crucial for refining these systems and ensuring their scalability and inclusivity in language instruction. Additionally, addressing the complexities of cultural context and learner variability will be key to the success of LLMs as effective agents in FLE.
7 Challenges and Future Directions
While we posit that LLMs have the potential to revolutionize FLE, realizing their full promise requires addressing key challenges. This section offers a concise overview of these challenges, followed by directions that could guide future research and deployment.
Ensuring Reliability and Mitigating Hallucinations.
LLMs may produce hallucinations (Huang et al., 2023) that can mislead learners and undermine pedagogical goals. This risk intensifies in high-stakes educational environments, where trust and correctness are paramount. Efforts to increase data quality (Long et al., 2024), combine model outputs with structured domain knowledge, and employ rigorous validation mechanisms (including human oversight) are crucial for minimizing such detrimental outcomes.
Addressing Bias and Ethical Considerations.
As LLMs inherit biases from their training data, these systems may produce culturally insensitive or unfair responses, potentially harming students from diverse linguistic and sociocultural backgrounds. Moreover, privacy concerns emerge when collecting and using learner data to personalize instruction. Robust governance frameworks, transparent documentation of data sources, and bias mitigation strategies (Borah & Mihalcea, 2024) can help ensure that LLMs serve as equitable tools in FLE.
Aligning With Pedagogical Principles.
LLMs excel at generating language but often lack pedagogical alignment, particularly for tasks requiring developmental sensitivity, motivation, or differentiated instruction. Their general-purpose nature means they do not inherently account for language acquisition theories or curricular standards (Razafinirina et al., 2024). To fully realize their potential, LLM-based tutors should be guided by robust pedagogical frameworks and co-designed with educators who understand learners’ cognitive and developmental needs.
8 Alternative Views
8.1 Task-Specific or Language-Specific Models as Better Alternatives
Some argue that specialized or language-specific models, including classical ML systems with carefully engineered features, can outperform general-purpose LLMs in narrowly defined tasks (e.g., phonetics or grammar drills (Fang et al., 2023)). By focusing on limited objectives, such models avoid the computational overhead and potential inaccuracies of LLMs, which aim to handle a broader range of inputs and contexts (Shen et al., 2024a).
Counterargument.
While specialized models may excel in isolated tasks, they lack the flexibility required for comprehensive FLE, which involves cultural nuances, conversations, and evolving learner needs. In contrast, LLMs can be fine-tuned for specific goals while still offering broader linguistic competence (Song et al., 2024a). Additionally, relying on multiple specialized models can be resource-intensive, whereas a well-configured LLM provides a unified framework that balances specialization and scalability.
8.2 Concerns About Over-Reliance on LLMs
Critics warn that over-reliance on LLMs may lead to problems such as generating misleading outputs (Nahar et al., 2024), reducing human interaction, and over-standardizing teaching methods. These issues could undermine the interpersonal and motivational aspects of language learning.
Counterargument.
These risks highlight the need for balanced integration rather than the replacement of human tutors. LLMs can complement educators by automating repetitive tasks, allowing teachers to focus on individualized support and motivation. Advances in AI safety, such as feedback loops (Tong et al., 2024) and human-in-the-loop systems (Wu et al., 2022), can help minimize inaccuracies (Ho et al., 2024). Additionally, the fine-tuning capabilities of LLMs ensure adaptability, supporting diverse and inclusive learning experiences (Lee et al., 2024b).
9 Conclusion
This position paper emphasizes the transformative potential of LLMs in FLE, positioning them as valuable tutors to complement traditional teaching methods. Through their roles as data enhancers, task predictors, and agents, LLMs can provide adaptive learning experiences across the core skills of listening, speaking, reading, and writing. This paper encourages interdisciplinary exploration to FLE.
Impact Statement
This paper positions LLMs as promising tutors in FLE. By leveraging LLMs, we aim to enhance inclusivity, personalization, and overall accessibility. However, potential risks—such as bias, privacy issues, and over-reliance on AI tutoring—necessitate careful oversight, which have been discussed in the paper. In particular, ensuring equitable data representation, model usage transparency, and human agency preservation remains critical. While careful integration with human expertise is essential, the potential for LLMs to revolutionize FLE is significant, paving the way for more inclusive, efficient, and engaging educational practices. We encourage continuing dialogue and interdisciplinary collaboration to responsibly integrate LLMs into educational ecosystems and mitigate potential harms.
Appendix A Literature Review
We provide an overview of LLM-centric research of FLE presented in Figure 4.
LLMs for Foreign Language Education
LLM as Data Enhancer (Section 4) Data Creation PFQS (Li & Zhang, 2024), MCQG-SRefine (Yao et al., 2024), (Lee et al., 2024c), BF-TC (Liu et al., 2024d), MathVC (Yue et al., 2024), EduAgent (Xu et al., 2024), Generative Students (Lu & Wang, 2024) Data Reformation Book2Dial (Wang et al., 2024b), Slide2Lecture (Zhang-Li et al., 2024), WikiDomains (Asthana et al., 2024), (Freyer et al., 2024), (Day et al., 2025), Anthropological Prompting (AlKhamissi et al., 2024), (Liu et al., 2024a) Data Annotation EXCGEC (Ye et al., 2024), EDEN (Siyan et al., 2024), MEGAnno+ (Kim et al., 2024), PDFChatAnnotator (Tang et al., 2024), Coannotating (Li et al., 2023) LLM as Task Predictor (Section 5) Discriminative TOREE (Zhuang et al., 2024), LLM-KT (Zhan et al., 2024), CLST (Jung et al., 2024), Diallogue-KT (Scarlatos & Lan, 2024), (Neshaei et al., 2024), (Mizumoto & Eguchi, 2023), (Seßler et al., 2024), (Syamkumar et al., 2024) Generative EXPECT (Fei et al., 2023), GEE (Song et al., 2024b), EXCGEC (Ye et al., 2024) FELT (Borges et al., 2024), LLM-as-a-tutor (Han et al., 2024), (Stamper et al., 2024) SocraticLM (Liu et al., 2024b), (Favero et al., 2024), (Xiao et al., 2024) Mixed FABRIC (Han et al., 2023), (Stahl et al., 2024b), ReaLMistake (Kamoi et al., 2024), (Lu et al., 2024) LLM-empowered Agent (Section 6) Abilities FOKE (Hu & Wang, 2024), KnowEdit (Zhang et al., 2024a), (Zhang et al., 2024a), PedCoT (Jiang et al., 2024b), LHP (Sonkar et al., 2024), (Razafinirina et al., 2024), LessonPlanner (Fan et al., 2024), MemoryBank (Zhong et al., 2024), (Hu et al., 2024), vNMF (Jiang et al., 2024a), ChatTutor (Chen et al., 2024), (Zheng et al., 2024) Applications SimClass (Zhang et al., 2024b), MathVC (Yue et al., 2024), BIPED (Kwon et al., 2024), LLM-as-a-tutor (Han et al., 2024), KORLINGS (Lee et al., 2024a), (Liu et al., 2024d)
Figure 4: An overview of LLM-centric research of FLE.
Appendix B Four Phases of Research Roadmap
Stage 1: Rule-based Models (1960s–1990s).
Early solutions relied on handcrafted linguistic rules to process language in tightly constrained scenarios (Grosan et al., 2011; C Angelides & Garcia, 1993). Classical platforms like PLATO (Hart, 1981) and Systran (Toma, 1977) operated effectively for highly structured tasks (e.g., grammar drills) but struggled with complex, context-dependent interactions.
Stage 2: Statistical Models (1990s–2010s).
With the increased availability of digitized corpora, methods such as the early version of Google Translate (Och, 2006) and Dragon NaturallySpeaking (Blair, 1997) pioneered statistical pattern mining. These approaches leveraged large datasets to infer linguistic rules probabilistically, improving scalability yet still lacking deeper semantic understanding.
Stage 3: Neural Models (2010s–2020s).
The advent of deep learning architectures (e.g., RNNs (Yu et al., 2019) and Transformers (Vaswani, 2017)) enabled more robust context modeling, sparking transformative applications like Grammarly (Fitria, 2021) and Duolingo (Vesselinov & Grego, 2012). These systems offered enhanced personalization and feedback, significantly augmenting learners’ writing and reading comprehension.
Stage 4: Large Language Models (2020s–Present).
Today’s LLMs (e.g., ChatGPT (Achiam et al., 2023)) combine massive pre-training corpora with generative capabilities, achieving impressive results in multi-turn dialogue, individualized scaffolding, and multimodal integration. Tools such as Khanmigo (Anand, 2023) demonstrate LLMs’ potential for real-time conversational practice, dynamic content creation, and inclusive educational support at scale.

### v2

@@
-Position: LLMs Can be Good Tutors in Foreign Language Education
+Position: LLMs Can be Good Tutors in English Education
 Abstract
-While recent efforts have begun integrating large language models (LLMs) into foreign language education (FLE), they often rely on traditional approaches to learning tasks without fully embracing educational methodologies, thus lacking adaptability to language learning.
-To address this gap, we argue that LLMs have the potential to serve as effective tutors in FLE.
+While recent efforts have begun integrating large language models (LLMs) into English education, they often rely on traditional approaches to learning tasks without fully embracing educational methodologies, thus lacking adaptability to language learning.
+To address this gap, we argue that LLMs have the potential to serve as effective tutors in English Education.
 Specifically, LLMs can play three critical roles: (1) as data enhancers, improving the creation of learning materials or serving as student simulations; (2) as task predictors, serving as learner assessment or optimizing learning pathway; and (3) as agents, enabling personalized and inclusive education.
-We encourage interdisciplinary research to explore these roles, fostering innovation while addressing challenges and risks, ultimately advancing FLE through the thoughtful integration of LLMs.
-Machine Learning, ICML
+We encourage interdisciplinary research to explore these roles, fostering innovation while addressing challenges and risks, ultimately advancing English Education through the thoughtful integration of LLMs.
+Position: LLMs Can be Good Tutors in English Education
+Jingheng Ye1,2, Shen Wang1, Deqing Zou2, Yibo Yan1,3, Kun Wang1, Hai-Tao Zheng2,4thanks: Corresponding Author., Ruitong Liu2, Zenglin Xu5, Irwin King6, Philip S.
+Yu7, Qingsong Wen [^1] footnotemark: 1 1Squirrel Ai Learning, 2Tsinghua University, 3HKUST (Guangzhou), 4Peng Cheng Laboratory, 5Fudan University, 6The Chinese University of Hong Kong, 7University of Illinois at Chicago yejh22@mails.tsinghua.edu.cn
 1 Introduction
-Foreign Language Education (FLE) has long been a cornerstone of global education and a critical component of K-12 curricula, equipping students with the linguistic and cultural competencies necessary for an interconnected world.
-However, traditional FLE methods often fall short in addressing the diverse needs of learners (Hou, 2020).
-Challenges such as limited personalization, scalability constraints, and the lack of real-time feedback are particularly pronounced in large classroom settings (Ehrenberg et al., 2001), where individual attention is scarce, leaving many students disengaged or struggling to keep pace.
-Addressing these shortcomings requires innovative approaches that not only enhance the quality of instruction but also adapt to the unique learning trajectories of students (Eaton, 2010).
-In parallel, the advent of LLMs has opened new possibilities for transforming FLE (Caines et al., 2023).
-These models, powered by large-scale pre-training, exhibit remarkable capabilities in natural language understanding and generation, making them promising candidates for roles traditionally filled by human tutors.
-By leveraging LLMs as AI tutors, we may overcome many of the inherent limitations of conventional teaching methods, offering scalable, interactive, and personalized learning experiences (Chen et al., 2024; Schmucker et al., 2024).
-Therefore, this position paper argues that LLMs can be effective tutors in foreign language education, complementing human expertise and addressing key limitations of traditional methods.
-Figure 1: Involved disciplines of LLM for FLE.
-Figure 2: Overview of roles of LLMs in foreign language education. An overview of related literature is provided in Appendix A.
-As shown in Figure 1, FLE intersects with multiple disciplines, each of which underscores the potential of LLMs to revolutionize this domain.
-From the perspective of (1) computer science, advancements in machine learning and NLP have enabled LLMs to process and generate human-like language at an unprecedented scale; (2) linguistics (Radford et al., 2009) contributes a deeper understanding of grammar, phonetics, and semantics, allowing LLMs to generate accurate and understandable language outputs; (3) education provides the foundation for designing effective pedagogical strategies, ensuring that LLMs can deliver personalized, engaging, and developmentally appropriate learning experiences; and finally, (4) psycholinguistics (Steinberg & Sciarini, 2013) bridges the gap between language acquisition and cognitive processes, enabling LLMs to optimize learner interactions by adapting to individual needs and fostering meaningful engagement.
-Together, these disciplines position LLMs as uniquely capable of addressing the multifaceted challenges of foreign language education.
-Moreover, FLE encompasses four core skills: listening, speaking, reading, and writing, each of which can be significantly enhanced by LLMs.
-For listening, LLMs can generate diverse audio materials (Ghosal et al., 2023) and facilitate interactive voice-based exercises, helping learners improve their ability to discern pronunciation, intonation, and contextual meaning.
-In speaking, LLMs can simulate realistic conversations (Siyan et al., 2024), provide pronunciation feedback, and scaffold learners’ oral communication skills through iterative practice.
-For reading, LLMs can curate leveled texts, generate comprehension questions (Samuel et al., 2024), and engage learners in discussions that deepen their understanding of written content.
-Finally, in writing, LLMs can offer real-time grammar, syntax, and style feedback while assisting with idea generation and iterative revisions (Stahl et al., 2024a).
+English Education has long been a cornerstone of global education and a critical component of K-12 curricula, equipping students with the linguistic and cultural competencies necessary for an interconnected world Alhusaiyan (2025); Katinskaia (2025).
+However, traditional English teaching methods often fall short in addressing the diverse needs of learners Hou (2020).
+Challenges such as limited personalization, scalability constraints, and the lack of real-time feedback are particularly pronounced in large classroom settings Ehrenberg et al.
+(2001).
+Addressing these shortcomings requires innovative approaches that not only enhance the quality of instruction but also adapt to the unique learning trajectories of students Eaton (2010).
+Recently, LLMs have opened new possibilities for transforming English Education Caines et al.
+(2023).
+LLMs exhibit remarkable natural language understanding and generation capabilities, making them promising candidates for roles traditionally filled by human tutors.
+Leveraging LLMs as AI tutors can overcome the inherent limitations of conventional teaching methods, offering scalable, interactive, and personalized learning experiences Chen et al.
+(2024); Schmucker et al.
+(2024).
+Therefore, this position paper argues that LLMs can be effective tutors in English education, complementing human expertise and addressing key limitations of traditional methods.
+Figure 1: Involved disciplines of LLM for English Edu.
+Figure 2: Overview of three roles of LLMs in English education. An overview of related literature is provided in Appendix A.
+As shown in Figure 1, English Education intersects with multiple disciplines, each of which underscores the potential of LLMs to revolutionize this domain.
+From the perspective of (1) computer science, advancements in machine learning and NLP have enabled LLMs to process and generate human-like language at an unprecedented scale; (2) linguistics Radford et al.
+(2009) contributes a deeper understanding of grammar, phonetics, and semantics, allowing LLMs to generate accurate and understandable language outputs; (3) education provides the foundation for designing effective pedagogical strategies, ensuring that LLMs can deliver personalized, engaging, and developmentally appropriate learning experiences; and finally, (4) psycholinguistics Steinberg and Sciarini (2013) bridges the gap between language acquisition and cognitive processes, enabling LLMs to optimize learner interactions by adapting to individual needs and fostering meaningful engagement.
+Together, these disciplines position LLMs as uniquely capable of addressing the multifaceted challenges of English education.
+Moreover, English Education encompasses four core skills: listening, speaking, reading, and writing, each of which can be significantly enhanced by LLMs.
+For listening, LLMs can generate diverse audio materials Ghosal et al.
+(2023) and facilitate interactive voice-based exercises, helping learners improve their ability to discern pronunciation, intonation, and contextual meaning.
+In speaking, LLMs can simulate realistic conversations Siyan et al.
+(2024), provide pronunciation feedback, and scaffold learners’ oral communication skills through iterative practice.
+For reading, LLMs can curate leveled texts, generate comprehension questions Samuel et al.
+(2024), and engage learners in discussions that deepen their understanding of written content.
+Finally, in writing, LLMs can offer real-time grammar, syntax, and style feedback while assisting with idea generation and iterative revisions Stahl et al.
+(2024a).
 By addressing these core skills holistically, LLMs have the potential to deliver a comprehensive and adaptive learning experience.
-Despite these opportunities, the deployment of LLMs in FLE must be approached carefully, ensuring that their integration complements rather than replaces human tutors (Jeon & Lee, 2023).
+Despite these opportunities, the deployment of LLMs in English Education must be approached carefully, ensuring that their integration complements rather than replaces human tutors Jeon and Lee (2023).
 As illustrated in Figure 2, this paper explores three critical roles of LLMs in this context: their function as data enhancers (Section 4) to optimize learning materials, their capacity as task predictors (Section 5) to tailor educational solutions, and their potential as agents (Section 6) that deliver interactive and adaptive language instruction.
-By examining these roles, we aim to demonstrate how LLMs can address the limitations of traditional FLE methods while advancing our understanding of intelligent tutoring systems.
-Additionally, we discuss potential challenges, ethical considerations, and future directions (Section 7) for integrating LLMs into FLE, offering a technical guideline for researchers and educators to harness their transformative potential.
-We also describe the paradigm shift of leveraging AI for FLE, starting from the last century, as one of our contributions in Section 3.
-Figure 3: Roadmap of Foreign Language Education.
+By examining these roles, we aim to demonstrate how LLMs can address the limitations of traditional English teaching methods while advancing our understanding of intelligent tutoring systems.
+Additionally, we discuss potential challenges (Section 7) and future directions (Appendix B) for integrating LLMs into English Education, offering a technical guideline for researchers and educators to harness their transformative potential.
+We also describe the paradigm shift of leveraging AI for English Education, starting from the last century, as one of our contributions in Section 3.
 2 Background
-2.1 Foreign Language Education
-FLE (Watzke, 2003; Rashov, 2024) has long been a cornerstone of global communication and cultural exchange.
-Traditional methods often emphasize grammar rules, vocabulary memorization, and repetitive practice, supplemented by limited opportunities for real-world application.
-Such approaches are often constrained by the availability of skilled teachers, the diversity of learners’ needs, and the lack of personalized feedback (Williams et al., 2004).
-To bridge the gaps, many technologies for FLE have been proposed (Alhusaiyan, 2024), focusing on solving specific tasks instead of describing the whole picture of foreign language tutoring.
-While intelligent language tutoring systems (Schwind, 1990; C Angelides & Garcia, 1993) have the potential to create adaptive environments, attention to this field is relatively less compared to other subjects like science and mathematics.
-One key reason lies in the inherent complexity of language as an ill-defined domain (Schmidt & Strasser, 2022), posing a great challenge in establishing a valid automatic analysis of learner languages due to the vast variability and unpredictability of human language.
-2.2 Large Language Models
-The potential of LLMs in education (Alhafni et al., 2024), particularly in FLE (Gao et al., 2024; Karataş et al., 2024), is immense.
-Benefiting from large-scale pre-training on extensive corpora, LLMs have demonstrated emergent abilities including (1) in-context learning (Dong et al., 2022), which allows the model to adapt to new tasks and provide contextually relevant responses based on a few examples provided during the interaction; (2) instruction following (Zeng et al., 2024), which enables the model to process and execute complex user instructions with high accuracy; and (3) reasoning and planning (Huang et al., 2024b), which allows the model to generate coherent, structured, and context-aware outputs, even for tasks that require multi-step thinking.
-However, these fundamental capabilities, while impressive, are not sufficient to fully meet the unique demands of FLE.
-Teaching a foreign language requires more than generating grammatically correct sentences or providing accurate translations; it demands a nuanced understanding of pedagogy, learner psychology, and cultural context.
+2.1 English Education
+Traditional English Education methods often emphasize grammar rules, vocabulary memorization, and repetitive practice, supplemented by limited opportunities for real-world application Watzke (2003).
+Such approaches are often constrained by the availability of skilled teachers, the diversity of learners’ needs, and the lack of personalized feedback Williams et al.
+(2004).
+Recently, many technologies for English Education have been proposed Alhusaiyan (2024), focusing on solving specific tasks instead of describing the whole picture of English tutoring.
+While intelligent language tutoring systems have the potential to create adaptive environments, attention to this field is relatively less compared to other subjects like science Shao et al.
+(2025) and mathematics Ahn et al.
+(2024).
+One key reason lies in the inherent complexity of language as an ill-defined domain Schmidt and Strasser (2022), posing a great challenge in establishing a valid automatic analysis of learner languages due to the vast variability and unpredictability of human language.
+2.2 Large Language Models for Education
+The potential of LLMs in education Alhafni et al.
+(2024), particularly in English Education Gao et al.
+(2024); Karataş et al.
+(2024); Cherednichenko et al.
+(2024), is immense.
+Benefiting from large-scale pre-training on extensive corpora, LLMs have demonstrated emergent abilities including (1) in-context learning Dong et al.
+(2022), which allows the model to adapt to new tasks and provide contextually relevant responses based on a few examples provided during the interaction; (2) instruction following Zeng et al.
+(2024), which enables the model to process and execute complex user instructions with high accuracy; and (3) reasoning and planning Huang et al.
+(2024b), which allows the model to generate coherent, structured, and context-aware outputs, even for tasks that require multi-step thinking.
+However, these fundamental capabilities, while impressive, are insufficient to fully meet the unique demands of English Education.
+Teaching English requires more than generating grammatically correct sentences or providing accurate translations; it demands a nuanced understanding of pedagogy, learner psychology, and cultural context.
 Maurya et al.
@@
 These dimensions can be broadly categorized into two groups.
-(1) Problem-solving abilities assess the technical capabilities of LLMs to perform tasks relevant to FLE.
+(1) Problem-solving abilities assess the technical capabilities of LLMs to perform tasks relevant to English Education.
 (2) Pedagogical alignment abilities evaluate how well the LLM aligns with effective teaching and learning principles.
 Pedagogical alignment includes the model’s ability to adapt to the learner’s proficiency level, provide scaffolded feedback, foster engagement, and maintain motivation.
-While LLMs can give direct answers, their ability to replicate these nuanced teaching strategies remains a challenge (Wang et al., 2024a).
-Additionally, pedagogical alignment requires sensitivity to cultural and contextual factors, ensuring that the learning experience is relevant and meaningful to the student.
+While LLMs can give direct answers, their ability to replicate these nuanced teaching strategies remains a challenge Wang et al.
+(2024a).
 3 Paradigm Shift
-The development of AI models for FLE can be broadly categorized into four successive generations as shown in Figure 3: (1) rule-based models, (2) statistical models, (3) neural models, and (4) large language models.
-We leave the detailed description in Appendix B.
+The development of AI models for English Education can be broadly categorized into four successive generations as shown in Figure 3: (1) rule-based models, (2) statistical models, (3) neural models, and (4) large language models.
+Stage 1: Rule-based Models (1960s–1990s).
+Early solutions relied on handcrafted linguistic rules to process language in tightly constrained scenarios Grosan et al.
+(2011); C Angelides and Garcia (1993).
+Classical platforms like PLATO Hart (1981) and Systran Toma (1977) operated effectively for highly structured tasks (e.g., grammar drills) but struggled with complex, context-dependent interactions.
+Stage 2: Statistical Models (1990s–2010s).
+With the increased availability of digitized corpora, methods such as the early version of Google Translate Och (2006) and Dragon NaturallySpeaking Blair (1997) pioneered statistical pattern mining.
+These approaches leveraged large datasets to infer linguistic rules and conduct specific tasks probabilistically, improving scalability yet still lacking deeper semantic understanding.
+Stage 3: Neural Models (2010s–2020s).
+The advent of deep learning architectures (e.g., RNNs Yu et al.
+(2019) and Transformers Vaswani (2017)) enabled more robust context modeling, sparking transformative applications like Grammarly Fitria (2021) and Duolingo Vesselinov and Grego (2012).
+These systems offered enhanced personalization and feedback, significantly augmenting learners’ writing and reading comprehension.
+Stage 4: Large Language Models (2020s–Present).
+Nowadays, various LLMs (e.g., ChatGPT Achiam et al.
+(2023)) combine massive pre-training with instruction tuning, achieving impressive results in multi-turn dialogue, individualized scaffolding, and multimodal integration.
+Tools such as Khanmigo Anand (2023) demonstrate LLMs’ potential for real-time conversational practice, dynamic content creation, and inclusive educational support at scale.
 Our position.
 We foresee next-generation LLMs with deeper alignment to pedagogical principles and stronger guardrails to mitigate misinformation and bias.
-Future models may integrate multimodal data (e.g., text, image, video, speech) and employ meta-learning to adapt in real time to diverse learner profiles.
-These improvements will reinforce the position that LLMs can evolve into even more effective tutors for FLE.
+Future models may integrate multimodal data (e.g., text, image, video, speech) to adapt to diverse learner profiles in real time.
+These improvements will reinforce the position that LLMs can evolve into more effective tutors for English Education.
+Figure 3: Roadmap of English Education.
 4 LLMs as Data Enhancers
-Education is a high-stake area where any form of LLMs’ hallucination could cause devastating harm to humans’ cognition activities (Ho et al., 2024).
-One of the hallucination causes is from data (Huang et al., 2023).
-Therefore, high-quality and diverse data resources (Long et al., 2024) are critical to ensuring the reliability of incorporating LLMs into FLE.
+Education is a high-stake area where any hallucination could cause devastating harm to humans’ cognition activities Ho et al.
+(2024).
+One of the hallucination causes is from data Huang et al.
+(2023).
+Therefore, high-quality and diverse data resources Long et al.
+(2024) are critical to ensuring the reliability of incorporating LLMs into English Education.
 The 1) creation, 2) reformation, and 3) annotation of educational materials are crucial to delivering effective and engaging teaching.
-Traditional resource development methods often lack the scalability, adaptability, and personalization necessary to meet the diverse needs of learners (Feng et al., 2021; Shorten et al., 2021).
-In contrast, LLMs emerge as transformative tools capable of enhancing these processes (Wang et al., 2024c; Liu et al., 2024c).
-This section explores how LLMs serve as data enhancers in FLE.
+Traditional resource development methods often lack the scalability, adaptability, and personalization necessary to meet the diverse needs of learners Feng et al.
+(2021); Shorten et al.
+(2021).
+In contrast, LLMs emerge as transformative tools capable of enhancing these processes Wang et al.
+(2024c); Liu et al.
+(2024c).
+This section explores how LLMs serve as data enhancers in English Education.
 4.1 Data Creation
 Creating pedagogically sound and learner-specific data is a cornerstone of personalized learning.
-However, manually creating such resources is time-consuming and often fails to address the wide range of learner needs (Cochran et al., 2022).
-LLMs can revolutionize this process by generating tailored and diverse educational content or responses on demand (Zha et al., 2023; Cochran et al., 2023).
+However, manually creating such resources is time-consuming and often fails to address the wide range of learner needs Cochran et al.
+(2022).
+LLMs can revolutionize this process by generating tailored and diverse educational content or responses on demand Zha et al.
+(2023); Cochran et al.
+(2023).
 Educational Materials Generation.
 A primary use of LLMs in data creation is the generation of educational questions aligned with specific learning objectives.
-Due to their superior contextual understanding, classic rule-based approaches have largely been eclipsed by neural network-based techniques (Kurdi et al., 2020; Rathod et al., 2022; Mulla & Gharpure, 2023).
-LLMs can produce answer-aware (whose target answer is known) or answer-agnostic (whose answer is open) (Zhang et al., 2021), resulting in more nuanced exercises and assessments (Xiao et al., 2023).
+Due to their superior contextual understanding, classic rule-based approaches have largely been eclipsed by neural network-based techniques Kurdi et al.
+(2020); Rathod et al.
+(2022); Mulla and Gharpure (2023).
+LLMs can produce answer-aware (whose target answer is known) or answer-agnostic (whose answer is open) Zhang et al.
+(2021), resulting in more nuanced exercises and assessments Xiao et al.
+(2023).
 Student Simulation.
 Simulating the learner’s perspective is crucial for designing adaptive instructional materials.
-Traditional surveys and standardized tests often fail to capture the complexity of dynamic learner behaviors (Käser & Alexandron, 2024).
-In contrast, LLM-based approaches enable high-fidelity, context-aware student simulations (Liu et al., 2024d; Yue et al., 2024), generating synthetic learners who exhibit realistic mastery levels and evolving behaviors.
-For instance, Generative Students (Lu & Wang, 2024) create simulated learners with various competency levels, while EduAgent (Xu et al., 2024) integrates cognitive priors to model complex learning trajectories and behaviors better.
+Traditional surveys and standardized tests often fail to capture the complexity of dynamic learner behaviors Käser and Alexandron (2024).
+In contrast, LLM-based approaches enable high-fidelity, context-aware student simulations Liu et al.
+(2024d); Yue et al.
+(2024), generating synthetic learners who exhibit realistic mastery levels and evolving behaviors.
+For instance, Generative Students Lu and Wang (2024) create simulated learners with various competency levels, while EduAgent Xu et al.
+(2024) integrates cognitive priors to model complex learning trajectories and behaviors better.
 Discussion.
-While LLMs excel at generating educational content, current approaches mainly focus on question creation, leaving many areas of FLE underexplored.
-Essential tasks like generating culturally rich reading materials, context-dependent writing prompts, or dynamic comprehension exercises are still lacking in diversity and depth.
-Additionally, the student simulations created by LLMs often fail to reflect long-term learning trajectories or the intricacies of individual learning progress, leading to challenges in creating truly adaptive systems.
+While LLMs excel at generating educational content, current approaches mainly focus on question creation, leaving many areas of English Education underexplored.
+Essential tasks like generating culturally rich reading materials, context-dependent writing prompts, or dynamic comprehension exercises still lack diversity and depth.
+Additionally, the student simulations created by LLMs often fail to reflect long-term learning trajectories or the intricacies of individual learning progress.
 4.2 Data Reformation
@@
 Transforming existing materials into different forms can yield more comprehensive and immersive learning experiences.
-For example, Book2Dial(Wang et al., 2024b) generates teacher-student dialogues grounded in textbooks, keeping the content both relevant and informative.
-Their approach includes multi-turn question generation and answering(Kim et al., 2022), dialogue inpainting (Dai et al., 2022), and role-playing.
-Likewise, Slide2Lecture (Zhang-Li et al., 2024) automatically converts lecture slides into structured teaching agendas, enabling interactive follow-up and deeper learner engagement.
+For example, Book2DialWang et al.
+(2024b) generates teacher-student dialogues grounded in textbooks, keeping the content both relevant and informative.
+Their approach includes multi-turn question generation and answeringKim et al.
+(2022), dialogue inpainting Dai et al.
+(2022), and role-playing.
+Likewise, Slide2Lecture Zhang-Li et al.
+(2024) automatically converts lecture slides into structured teaching agendas, enabling interactive follow-up and deeper learner engagement.
 Simplification and Paraphrasing.
-Another vital application is simplifying or paraphrasing complex texts to specified readability levels (Huang et al., 2024a) without losing key concepts (Al-Thanyyan & Azmi, 2021).
-This is particularly beneficial in FLE settings, where language beginners often face advanced vocabulary and complex structures (Day et al., 2025).
-Recent advancements in controllable generation (Zhang et al., 2023) leverage model fine-tuning on curated datasets (Zeng et al., 2023) or decoding-time interventions (Liang et al., 2024), thereby allowing educators to specify text complexity, style, or tone.
+Another vital application is simplifying or paraphrasing complex texts to specified readability levels Huang et al.
+(2024a) without losing key concepts Al-Thanyyan and Azmi (2021).
+This is particularly beneficial in English Education settings, where language beginners often face advanced vocabulary and complex structures Day et al.
+(2025).
+Recent advancements in controllable generation Zhang et al.
+(2023) leverage model fine-tuning on curated datasets Zeng et al.
+(2023) or decoding-time interventions Liang et al.
+(2024), thereby allowing educators to specify text complexity, style, or tone.
 Cultural Context Adaptation.
-Beyond linguistic correctness, cultural nuance is pivotal in FLE (Byram, 1989, 2008).
-LLMs can facilitate this process by recontextualizing existing materials to reflect the cultural and social norms of the target language (Liu et al., 2024a).
-For instance, a short story originally set in an English-speaking environment may be adapted for French or Japanese learners by adjusting the characters’ names, idiomatic expressions, or social customs, while still preserving core instructional goals.
+Beyond linguistic correctness, cultural nuance is another crucial factor in English Education Byram (1989, 2008).
+LLMs can facilitate this process by recontextualizing existing materials to reflect the cultural and social norms of different areas Liu et al.
+(2024a); Adilazuarda et al.
+(2024); Kharchenko et al.
+(2024).
+For instance, a short story originally set in an English-speaking environment may be adapted for Japanese students by adjusting the characters’ names, idiomatic expressions, or social customs, while preserving core instructional goals.
 This cultural adaptation not only enhances learner engagement but also strengthens cross-cultural competencies.
 Discussion.
-While LLM-based data reformation can significantly enhance FLE, several gaps warrant attention.
-Most current studies prioritize textual forms or single-modal approaches, which may overlook valuable multimodal resources such as interactive video and audio-based content (Ghosal et al., 2023).
-Furthermore, cultural adaptation, although promising, remains underexplored in practical classroom scenarios, particularly for underrepresented personas and culturally sensitive topics.
+While LLM-based data reformation can significantly enhance English Education, several gaps warrant attention.
+Most current studies prioritize textual forms or single-modal approaches, which may overlook valuable multimodal resources such as interactive video and audio-based content Ghosal et al.
+(2023).
+Furthermore, cultural adaptation, although promising, remains underexplored in practical classroom scenarios, particularly for underrepresented persons and culturally sensitive topics.
 AlKhamissi et al.
-(2024) demonstrate how cultural misalignment can increase bias, suggesting Anthropological Prompting to strengthen cultural alignment.
+(2024) demonstrate how cultural misalignment can increase bias.
 However, robust empirical evaluations are still limited across diverse learners and linguistic backgrounds.
-These issues highlight the importance of interdisciplinary research, where educators, linguists, and technologists collaborate to refine data reformation techniques that broaden inclusivity and effectiveness in FLE.
 4.3 Data Annotation
 While Data Creation focuses on generating learner-specific data, it often prioritizes diversity and adaptability over precision.
-The approach is particularly useful for tasks with large label spaces (Ding et al., 2024).
+The approach is particularly useful for tasks with large label spaces Ding et al.
+(2024).
 In contrast, Data Annotation emphasizes producing high-quality, meticulously labeled data that is essential for tasks requiring accuracy and consistency.
-Unlike data creation, annotated data often undergoes rigorous validation to ensure its accuracy and relevancy (Artemova et al., 2024).
+Unlike data creation, annotated data often undergoes rigorous validation to ensure its accuracy and relevancy Artemova et al.
+(2024).
 Annotation Generation.
@@
 (2024) deploy GPT-4-Turbo for audio transcript annotations.
-However, inconsistencies across LLMs (Törnberg, 2024) remain a serious challenge, posing risks to educational reliability.
+However, inconsistencies across LLMs Törnberg (2024) remain a serious challenge, posing risks to educational reliability.
 Annotation Assessment.
-Although LLM-based annotation generation is efficient, it also raises critical issues of bias, calibration, and validity—particularly in low-resource language contexts (Bhat & Varma, 2023; Jadhav et al., 2024).
+Although LLM-based annotation is efficient, it also raises critical issues of bias, calibration, and validity, particularly in low-resource language contexts Bhat and Varma (2023); Jadhav et al.
+(2024).
 Automated or semi-automated evaluation strategies have emerged to address these quality concerns.
-For example, LLMs-as-Judges (Li et al., 2024a, b; Gu et al., 2024) reduce human overhead by automating evaluation, an approach increasingly explored in education-focused applications (Chiang et al., 2024; Zhou et al., 2024).
+For example, LLMs-as-Judges Li et al.
+(2024a, b); Gu et al.
+(2024) reduce human overhead by automating evaluation, an approach increasingly explored in education-focused applications Chiang et al.
+(2024); Zhou et al.
+(2024).
 However, purely automated frameworks can still propagate errors or bias.
@@
 Although LLMs provide efficient data annotation, the inconsistency across different models remains a critical concern, affecting the quality and reliability of annotated educational materials.
-These discrepancies hinder the creation of universally reliable FLE content, especially in diverse linguistic and cultural contexts.
+These discrepancies hinder the creation of universally reliable educational content, especially in diverse linguistic and cultural contexts.
 Additionally, automated annotations often lack the nuance needed for pedagogical applications, making it essential to involve human oversight in critical cases to mitigate errors or biases.
 Our position.
-We acknowledge the current limitations in LLM-based data creation, reformation, and annotation for FLE.
-However, we believe that with continued interdisciplinary collaboration, including input from educators, linguists, and technologists, these challenges can be addressed.
-Future advancements should focus on enhancing the accuracy and diversity of generated content, improving multimodal and culturally sensitive learning materials, and integrating more robust systems for human-LLM collaboration (Li et al., 2023; Wang et al., 2024e) in data annotation.
-This approach will ensure that LLMs can fully realize their potential as effective tutors in FLE.
+We acknowledge the current limitations in LLM-based data creation, reformation, and annotation for English Education.
+However, we believe that with continued interdisciplinary collaboration, these challenges can be addressed.
+Future advancements should focus on enhancing the accuracy and diversity of generated content, improving multi-modal and culturally sensitive learning materials, and integrating more robust systems for human-LLM collaboration Li et al.
+(2023); Wang et al.
+(2024e) in data annotation.
+This will ensure that LLMs can fully realize their potential as effective tutors in English Education.
 5 LLMs as Task Predictors
-Task-Based Language Learning (TBLL) (Nunan, 1989; Willis, 2021) as a methodological approach is one of the critical factors of modern foreign language teaching.
-LLMs have demonstrated remarkable capabilities in understanding and generating human language, making them well-suited for addressing numerous tasks in FLE.
+Task-Based Language Learning (TBLL) Nunan (1989); Willis (2021) as a methodological approach is one of the effective English Education methods.
+LLMs have demonstrated remarkable capabilities in understanding and generating human language, making them well-suited for addressing numerous tasks in English Education.
 These tasks can be broadly categorized into three types based on their nature and the role of LLMs: 1) Discriminative, 2) Generative, and 3) Mixed of the above two roles.
 5.1 Discriminative Task Predictors
-Discriminative tasks in FLE primarily involve classifying learner inputs or grading their future performance. Below are some applications that are still calling for improvements:
+Discriminative tasks in English Education primarily involve classifying learner inputs or grading their future performance. Below are some applications that are still calling for improvements:
 Automated Assessment.
-The task aims to automatically grade students’ assignments, including essay scoring (Seßler et al., 2024; Li & Liu, 2024; Syamkumar et al., 2024), short answer grading (Schneider et al., 2023; Henkel et al., 2024), and spoken language evaluation (Gao et al., 2023; Fu et al., 2024).
+The task aims to automatically grade students’ assignments, including essay scoring Seßler et al.
+(2024); Li and Liu (2024); Syamkumar et al.
+(2024), short answer grading Schneider et al.
+(2023); Henkel et al.
+(2024), and spoken language evaluation Gao et al.
+(2023); Fu et al.
+(2024).
 LLMs can process learners’ submissions to judge grammar, lexical diversity, coherence, and even spoken fluency, providing instant feedback.
-This scalability is particularly appealing for large classes, where human evaluators are often overwhelmed and unable to provide timely, personalized critique (Mizumoto & Eguchi, 2023).
+This scalability is particularly appealing for large classes, where human evaluators are often overwhelmed and unable to provide timely, personalized critique Mizumoto and Eguchi (2023).
 Knowledge Tracing.
-Given sequences of learning interactions in online learning systems, Knowledge Tracing identifies and tracks students’ evolving mastery of target skills (Shen et al., 2024b; Xu et al., 2023).
-LLM-based methods have been explored in cold-start scenarios (Zhan et al., 2024; Jung et al., 2024), offering strong generalization by inferring latent learner states from limited data.
+Given sequences of learning interactions in online learning systems, Knowledge Tracing identifies and tracks students’ evolving mastery of target skills Shen et al.
+(2024b); Xu et al.
+(2023).
+LLM-based methods of Knowledge Tracing have been explored in cold-start scenarios Zhan et al.
+(2024); Jung et al.
+(2024), offering strong generalization by inferring latent learner states from limited data.
 These approaches can support adaptive learning pathways, giving personalized recommendations based on predicted performance and knowledge gaps.
@@
 Despite their promise in automating and personalizing these discriminative tasks, LLMs still grapple with notable limitations that hinder their utility as robust tutoring tools.
-First, misalignment of assessment with expert instructors poses risks: machine-generated scores may deviate from established rubrics or neglect qualitative nuances, leading to potential discrepancies in grading quality (Kundu & Barbosa, 2024).
-Second, the lack of empathy compounds this issue, as assessments devoid of human judgment risk discouraging learners or overlooking subtle motivational factors (Sharma et al., 2024).
-Knowledge tracing approaches, while promising in cold-start scenarios, struggle with capturing the complexity of long-term learning trajectories and deeper cognitive processes (Cho et al., 2024).
+First, misalignment of assessment with expert instructors poses risks: machine-generated scores may deviate from established rubrics or neglect qualitative nuances, leading to potential discrepancies in grading quality Kundu and Barbosa (2024).
+Second, the lack of empathy compounds this issue, as assessments devoid of human judgment risk discouraging learners or overlooking subtle motivational factors Sharma et al.
+(2024).
+Knowledge tracing approaches, while promising in cold-start scenarios, struggle with capturing the complexity of long-term learning trajectories and deeper cognitive processes Cho et al.
+(2024).
 These concerns point to the need for more transparent and human-centered methods in utilizing LLMs for assessment.
 5.2 Generative Task Predictors
-Generative tasks involve producing new content or responses. LLMs are known to be adept in these tasks due to their natural language generation capabilities.
+Generative tasks involve producing new content or responses. LLMs are known to be adept at these tasks due to their generation capabilities.
 Grammatical Error Correction and Explanation.
-In foreign language writing, errors often reveal learners’ gaps in grammar and vocabulary (Hyland & Hyland, 2006).
-LLMs can detect and correct these errors (Bryant et al., 2023; Ye et al., 2023), offering concise explanations (Ye et al., 2024) that reinforce language rules.
-By streamlining error detection and pedagogically framing corrections, learners deepen their linguistic understanding.
+In English writing, errors often reveal learners’ gaps in grammar and vocabulary Hyland and Hyland (2006).
+LLMs can detect and correct these errors Bryant et al.
+(2023); Ye et al.
+(2023), offering concise explanations Ye et al.
+(2024) that reinforce language rules.
+By streamlining error detection and corrections, learners deepen their linguistic understanding.
 Feedback Generation.
-Quizzes and exercises remain vital in FLE for practice and targeted remediation (Rashov, 2024).
-LLMs enhance this process by delivering prompt, personalized feedback that pinpoints strengths and addresses weaknesses (Borges et al., 2024).
-This scalability enables learners to self-regulate and refine their skills without relying solely on human graders (Stamper et al., 2024).
+Quizzes and exercises remain vital in English Education for practice and targeted remediation Rashov (2024).
+LLMs enhance this process by delivering prompt, personalized feedback that pinpoints strengths and addresses weaknesses Borges et al.
+(2024).
+This scalability enables learners to self-regulate and refine their skills without relying solely on human graders Stamper et al.
+(2024).
 Socratic Dialogue.
-Moving beyond straightforward Q&A, Socratic questioning promotes critical thinking and self-reflection (Paul & Elder, 2007).
-SocraticLM (Liu et al., 2024b), for example, aligns an LLM with open-ended, inquiry-based teaching principles, guiding learners through iterative exploration rather than prescriptive correction.
+Moving beyond straightforward Q&A, Socratic questioning promotes critical thinking and self-reflection Paul and Elder (2007).
+SocraticLM Liu et al.
+(2024b), for example, aligns an LLM with open-ended, inquiry-based teaching principles, guiding learners through iterative exploration rather than prescriptive correction.
 In theory, this fosters deeper conceptual understanding and active learner engagement.
 Discussion.
-Despite the promise of LLM-based generation in FLE, multiple uncertainties persist.
-Determining how to provide automatic feedback that genuinely maximizes learning outcomes is an ongoing challenge (Stamper et al., 2024), particularly given education’s risk-averse culture and high accountability standards (Xiao et al., 2024).
-Moreover, while LLMs like SocraticLM have demonstrated success in domains like mathematics, their applicability to FLE contexts has not been thoroughly validated (Liu et al., 2024b).
-As such, the design of strategies and follow-up queries remains an open question in ensuring that these systems track and respond to learners’ cognitive states effectively.
+Despite the promise of LLM-based generation in English Education, multiple uncertainties persist.
+Determining how to provide automatic feedback that genuinely maximizes learning outcomes is an ongoing challenge Stamper et al.
+(2024), particularly given education’s risk-averse culture and high accountability standards Xiao et al.
+(2024).
+Moreover, while LLMs like SocraticLM have demonstrated success in domains like mathematics, their applicability to English Education contexts has not been thoroughly validated Liu et al.
+(2024b).
+As such, the design of strategies and follow-up queries remains an open question in ensuring these systems track and respond to learners’ cognitive states.
 5.3 Mixed Task Predictors
@@
 Automated Assessment with Feedback.
-While discriminative systems for essay scoring and speech evaluation primarily focus on assigning grades, LLMs extend these capabilities by simultaneously generating formative feedback (Katuka et al., 2024; Stahl et al., 2024b).
-For example, an LLM can evaluate the coherence and lexical diversity of a written assignment, then offer specific revision strategies.
+While discriminative systems for automated essay scoring and speech evaluation primarily focus on assigning grades, LLMs extend these capabilities by simultaneously generating formative feedback Katuka et al.
+(2024); Stahl et al.
+(2024b).
+For example, an LLM can evaluate the coherence and lexical diversity of a written essay, then offer specific revision strategies.
 In speaking practice, it can measure fluency and pronunciation accuracy while suggesting drills to refine intonation or stress patterns.
@@
 Error Analysis.
-Error Analysis systematically uncovers and categorizes learners’ missteps, from syntactic lapses in writing to flawed pronunciations in speaking (James, 2013; Erdoğan, 2005).
-LLMs functioning in a mixed capacity can classify these errors and generate corrective guidance, providing revised sentences, clarifications of grammatical rules, or remediation exercises for identified weaknesses (Myles, 2002; Mashoor & Abdullah, 2020).
+Error Analysis systematically uncovers and categorizes learners’ missteps, from syntactic lapses in writing to flawed pronunciations in speaking James (2013); Erdoğan (2005).
+LLMs functioning in a mixed capacity can classify these errors and generate corrective guidance, providing revised sentences, clarifications of grammatical rules, or remediation exercises for identified weaknesses Myles (2002); Mashoor and Abdullah (2020).
 Such insight facilitates targeted interventions that enhance language proficiency across modalities, including reading and listening.
@@
 Mixed-task systems hold promise by combining assessment and feedback generation, but they face notable challenges.
-One major issue is the weak alignment between scoring mechanisms and the quality of feedback provided (Stahl et al., 2024b).
+One major issue is the weak alignment between scoring mechanisms and the quality of feedback provided Stahl et al.
+(2024b).
 For example, while essay scoring systems may deliver comprehensive evaluations, the feedback often lacks specificity, limiting its instructional value.
-Additionally, although error analysis has potential, the absence of standardized pedagogical benchmarks, especially in oral tasks, hampers the reliability and comparability of LLM-based tools (Leu Jr, 1982).
-Furthermore, by addressing errors across these modalities, LLMs can deliver a holistic and adaptive learning experience that fosters deeper language proficiency (Zhao et al., 2024).
+Additionally, although error analysis has potential, the absence of standardized pedagogical benchmarks, especially in oral tasks, hampers the reliability and comparability of LLM-based tools Leu Jr (1982).
 Our position.
-While LLMs offer scalable solutions for task prediction in FLE, their current limitations—such as misalignment with expert assessments, lack of empathy, and weak alignment between assessment and feedback—require ongoing refinement.
+While LLMs offer scalable solutions for task prediction in English Education, their current limitations—such as misalignment with expert assessments, lack of empathy, and weak alignment between assessment and feedback—require ongoing refinement.
 Future research should focus on improving model transparency, enhancing the cultural and emotional sensitivity of LLMs, and refining task predictors to better reflect long-term learning trajectories and learner motivation.
-Additionally, developing standardized pedagogical benchmarks for error analysis, especially in oral tasks, will help ensure the consistency and reliability of LLM-generated feedback.
+Additionally, developing standardized pedagogical benchmarks for error analysis will help ensure the consistency and reliability of LLM-generated feedback.
 6 LLM-empowered Agent
-In this section, we delve into the potential of LLMs as intelligent agents in FLE.
+In this section, we delve into the potential of LLMs as intelligent tutoring agents in English Education.
 LLMs can act as catalysts for personalized learning, addressing the long-standing scalability, adaptability, and inclusivity challenges in traditional teaching paradigms.
@@
 Knowledge Integration.
-LLMs excel at merging structured educational knowledge graphs (Abu-Rasheed et al., 2024; Hu & Wang, 2024) with unstructured textual data (Li et al., 2024c; Modran et al., 2024), providing rich, contextualized information on linguistic constructs and cultural nuances.
-Their ability to perform real-time knowledge editing (Wang et al., 2024d; Zhang et al., 2024a) ensures learners receive content aligned with evolving language usage, addressing the inherent limitations of static materials.
+LLMs excel at merging structured educational knowledge graphs Abu-Rasheed et al.
+(2024); Hu and Wang (2024) with unstructured textual data Li et al.
+(2024c); Modran et al.
+(2024), providing rich, contextualized information on linguistic constructs and cultural nuances.
+Their ability to perform real-time knowledge editing Wang et al.
+(2024d); Zhang et al.
+(2024a) ensures learners receive content aligned with evolving language usage, addressing the inherent limitations of static materials.
 Pedagogical Alignment.
-LLMs require embedding with pedagogical principles to facilitate genuine learning experiences (Carroll, 1965; Taneja, 1995).
-Recent work incorporates theoretical frameworks, such as Bloom’s taxonomy (Bloom et al., 1956), to guide LLMs in systematically addressing different cognitive levels (Jiang et al., 2024b).
-Approaches like Pedagogical Chain of Thought (Jiang et al., 2024b) and preference learning (Sonkar et al., 2024; Rafailov et al., 2024) focus on aligning model responses with educational objectives.
+LLMs require embedding with pedagogical principles to facilitate genuine learning experiences Carroll (1965); Taneja (1995).
+Recent work incorporates theoretical frameworks, such as Bloom’s taxonomy Bloom et al.
+(1956), to guide LLMs in systematically addressing different cognitive levels Jiang et al.
+(2024b).
+Approaches like Pedagogical Chain of Thought Jiang et al.
+(2024b) and preference learning Sonkar et al.
+(2024); Rafailov et al.
+(2024) focus on aligning model responses with educational objectives.
 Planning.
-By assisting in crafting teaching objectives and lesson designs, LLMs can handle complex tasks such as differentiated instruction (Hu et al., 2024).
-LessonPlanner (Fan et al., 2024) has been proposed to assist novice teachers in preparing lesson plans, with expert interviews confirming its effectiveness.
+By assisting in crafting teaching objectives and lesson designs, LLMs can handle complex tasks such as differentiated instruction Hu et al.
+(2024).
+LessonPlanner Fan et al.
+(2024) has been proposed to assist novice teachers in preparing lesson plans, with expert interviews confirming its effectiveness.
 Zheng et al.
@@
 Memory.
-Effective tutoring systems track learner histories and tailor subsequent interactions accordingly (Jiang et al., 2024a; Chen et al., 2024).
+Effective tutoring systems track learner histories and tailor subsequent interactions accordingly Jiang et al.
+(2024a); Chen et al.
+(2024).
 When serving as memory-augmented agents, LLMs can retain individualized data—such as repeated grammar mistakes or overlooked vocabulary—thereby improving continuity and enabling consistent scaffolding of future learning tasks.
 Tool Using.
-Beyond textual interactions, LLM-based agents can integrate specialized tools to streamline the educational ecosystem, from cognitive diagnosis modules (Ma & Guo, 2019) to report generators (Zhou et al., 2025).
+Beyond textual interactions, LLM-based agents can integrate specialized tools to streamline the educational ecosystem, from cognitive diagnosis modules Ma and Guo (2019) to report generators Zhou et al.
+(2025).
 By orchestrating these resources, LLMs seamlessly unify diverse utilities under a single interface, enhancing learner experience and instructional efficiency.
-Discussion.
-Existing research often overlooks the interplay among listening, speaking, reading, and writing in real-world language learning (Caines et al., 2023; Shetye, 2024).
-Most systems concentrate on text-oriented features, lacking robust benchmarks and methodologies to evaluate integrated, multimodal interactions.
-Recent work on pedagogical alignment (Razafinirina et al., 2024) largely addresses textual data, leaving out real-time speaking and listening dynamics that demand complex, rapid feedback.
-Similarly, while memory modules can track repeated written errors, their effectiveness in monitoring and improving learning efficacy remains underexplored.
 6.2 Applications
-Although still in its early stages, LLM-empowered agents have already started to show promising applications in FLE.
+Although still in its early stages, LLM-empowered agents have already started to show promising applications in English Education.
 Classroom Simulation.
-Classroom simulation leverages LLM-empowered agents to recreate complex, interactive learning settings without the logistical hurdles of organizing physical classrooms (Zhang et al., 2024b).
+Classroom simulation leverages LLM-empowered agents to recreate complex, interactive learning settings without the logistical hurdles of organizing physical classrooms Zhang et al.
+(2024b).
 By simulating virtual students and tutors, researchers can study pedagogical strategies at scale, generate diverse learner interactions, and refine teaching techniques.
-Moreover, this virtual data can be used to fine-tune LLMs for specific educational contexts and learner profiles (Liu et al., 2024b), offering a cost-effective and adaptable approach to language instruction.
+Moreover, this virtual data can be used to fine-tune LLMs for specific educational contexts and learner profiles Liu et al.
+(2024b), offering a cost-effective and adaptable approach to language instruction.
 Intelligent Tutoring System (ITS).
-LLM-based agents have demonstrated the capacity to provide dynamic, personalized tutoring experiences (Kwon et al., 2024), effectively identifying learner weaknesses through large-scale linguistic analysis (Caines et al., 2023).
+LLM-based agents have demonstrated the capacity to provide dynamic, personalized tutoring experiences Wang et al.
+(2025); Kwon et al.
+(2024), effectively identifying learner weaknesses through large-scale linguistic analysis Caines et al.
+(2023).
 This makes them promising for delivering individualized instruction at scale.
-Although current ITS applications in mathematics (Pal Chowdhury et al., 2024) and science (Stamper et al., 2024) have shown success, the extension to FLE requires nuanced handling of cultural and contextual elements, as well as the unpredictability of human language usage.
+Although current ITS applications in mathematics Pal Chowdhury et al.
+(2024) and science Stamper et al.
+(2024) have shown success, the extension to English Education requires nuanced handling of cultural and contextual elements, as well as the unpredictability of human language usage.
 Discussion.
 Despite the promise of these applications, critical challenges remain.
-Existing classroom simulation frameworks often lack standardized benchmarks for FLE, making it difficult to assess the efficacy and generalizability of developed systems (Zhang et al., 2024b).
+Existing classroom simulation frameworks often lack standardized benchmarks for English Education, making it difficult to assess the efficacy and generalizability of developed systems Zhang et al.
+(2024b).
 In addition, evaluating language-specific tutoring strategies, including real-time conversational practice and holistic skill integration, remains an underexplored frontier.
-Addressing these gaps requires new datasets and metrics centered on holistic skill development, as well as interdisciplinary collaboration.
+Addressing these gaps requires new datasets and metrics centered on holistic skill development and interdisciplinary collaboration.
 Our position.
-We argue that to overcome current limitations, future research should focus on the integration of multimodal learning tasks (Sonlu et al., 2024) and the development of standardized frameworks for evaluating FLE simulations.
+We argue that future research should focus on integrating multimodal learning tasks Sonlu et al.
+(2024) and developing standardized frameworks for evaluating English Education simulations.
 Moreover, LLMs should evolve beyond text-based capabilities to provide real-time, context-sensitive feedback, particularly in speaking and listening.
-Interdisciplinary collaboration and the creation of new datasets tailored to FLE are crucial for refining these systems and ensuring their scalability and inclusivity in language instruction.
-Additionally, addressing the complexities of cultural context and learner variability will be key to the success of LLMs as effective agents in FLE.
-7 Challenges and Future Directions
-While we posit that LLMs have the potential to revolutionize FLE, realizing their full promise requires addressing key challenges.
+Interdisciplinary collaboration and the creation of new datasets tailored to English Education are crucial for refining these systems and ensuring their scalability and inclusivity in language instruction.
+7 Challenges
+While we posit that LLMs have the potential to revolutionize English Education, realizing their full promise requires addressing key challenges.
 This section offers a concise overview of these challenges, followed by directions that could guide future research and deployment.
 Ensuring Reliability and Mitigating Hallucinations.
-LLMs may produce hallucinations (Huang et al., 2023) that can mislead learners and undermine pedagogical goals.
+LLMs may produce hallucinations Huang et al.
+(2023) that can mislead learners and undermine pedagogical goals.
 This risk intensifies in high-stakes educational environments, where trust and correctness are paramount.
-Efforts to increase data quality (Long et al., 2024), combine model outputs with structured domain knowledge, and employ rigorous validation mechanisms (including human oversight) are crucial for minimizing such detrimental outcomes.
+Future directions include enhancing data quality and diversity for training Long et al.
+(2024), developing techniques to integrate LLM outputs with structured domain knowledge and pedagogical rules, and employing rigorous automated and human-in-the-loop validation mechanisms to minimize such detrimental outcomes and improve the factual grounding of LLM-generated educational content.
 Addressing Bias and Ethical Considerations.
-As LLMs inherit biases from their training data, these systems may produce culturally insensitive or unfair responses, potentially harming students from diverse linguistic and sociocultural backgrounds.
-Moreover, privacy concerns emerge when collecting and using learner data to personalize instruction.
-Robust governance frameworks, transparent documentation of data sources, and bias mitigation strategies (Borah & Mihalcea, 2024) can help ensure that LLMs serve as equitable tools in FLE.
+As LLMs inherit biases from their training data, these systems may produce culturally insensitive, stereotypical, or unfair responses, potentially harming students from diverse linguistic and sociocultural backgrounds.
+Moreover, significant privacy concerns emerge when collecting and using learner data to personalize instruction, particularly for K-12 students.
+Future research must focus on developing robust governance frameworks, transparent documentation of data sources and model behaviors, and advanced bias detection and mitigation strategies Borah and Mihalcea (2024); He and Li (2024) to ensure that LLM-based tools for English Education are equitable, fair, and uphold stringent data protection standards.
 Aligning With Pedagogical Principles.
-LLMs excel at generating language but often lack pedagogical alignment, particularly for tasks requiring developmental sensitivity, motivation, or differentiated instruction.
-Their general-purpose nature means they do not inherently account for language acquisition theories or curricular standards (Razafinirina et al., 2024).
-To fully realize their potential, LLM-based tutors should be guided by robust pedagogical frameworks and co-designed with educators who understand learners’ cognitive and developmental needs.
-8 Alternative Views
-8.1 Task-Specific or Language-Specific Models as Better Alternatives
-Some argue that specialized or language-specific models, including classical ML systems with carefully engineered features, can outperform general-purpose LLMs in narrowly defined tasks (e.g., phonetics or grammar drills (Fang et al., 2023)).
-By focusing on limited objectives, such models avoid the computational overhead and potential inaccuracies of LLMs, which aim to handle a broader range of inputs and contexts (Shen et al., 2024a).
+LLMs excel at generating fluent language but often lack deep pedagogical alignment, particularly for tasks requiring developmental sensitivity, learner motivation strategies, or differentiated instruction tailored to individual learning needs.
+Their general-purpose nature means they do not inherently account for established language acquisition theories or specific curricular standards Razafinirina et al.
+(2024).
+A crucial future direction is the development of methodologies to better imbue LLMs with pedagogical intelligence.
+This includes co-designing LLM applications with educators, fine-tuning models on high-quality pedagogical interaction data, and creating architectures that can dynamically adapt to learners’ cognitive states and developmental needs in English language learning.
+8 Conclusion
+This paper emphasizes the transformative potential of LLMs in English Education, positioning them as valuable tutors to complement traditional teaching methods.
+Through their roles as data enhancers, task predictors, and agents, LLMs can provide adaptive learning experiences across the core skills of listening, speaking, reading, and writing.
+This paper encourages continuing dialogue and interdisciplinary collaboration to responsibly integrate LLMs into educational ecosystems.
+Limitations
+Emphasis on potential over practical implementation barriers.
+This paper primarily focuses on the potential of LLMs to serve as effective tutors in English Education, outlining beneficial roles as data enhancers, task predictors, and agents.
+While we acknowledge the existence of challenges (to be discussed in Appendix 7), a limitation of this position is that the main arguments may not fully capture the considerable practical, socio-economic, and infrastructural hurdles that could impede the equitable and effective implementation of these LLM roles across diverse global educational contexts and resource settings.
+Generalizability and contextual adaptation of proposed roles.
+We propose three broad roles for LLMs in English Education.
+However, this paper does not provide an exhaustive analysis of how the efficacy and suitability of LLMs in these roles might vary significantly across different target languages (especially low-resource languages), specific learner demographics (e.g., preschoolers vs.
+K-12 vs.
+adult learners, learners with disabilities), diverse cultural contexts, or varying pedagogical philosophies.
+The general framework presented may require substantial adaptation and further research to be effectively applied in specific English Education scenarios.
+Nuances of human-LLM pedagogical interaction.
+While advocating for LLMs as tutors that can complement human expertise, this position paper does not delve deeply into the complex dynamics of the pedagogical interactions between learners, LLM-based tutors, human educators, and parents.
+Critical aspects such as optimizing the collaborative model, designing effective training for educators to leverage LLMs, mitigating risks of learner over-reliance, and ensuring that LLM interactions foster deep learning rather than superficial engagement are multifaceted issues that warrant more extensive investigation than afforded by the scope of this paper.
+Acknowledgements
+This research is supported in part by NSF under grants III-2106758, and POSE-2346158.
+This research is also supported by National Natural Science Foundation of China (Grant No.62276154); Research Center for Computer Network (Shenzhen) Ministry of Education, the Natural Science Foundation of Guangdong Province (Grant No.2023A1515012914 and 440300241033100801770); Basic Research Fund of Shenzhen City (Grant No.JCYJ20210324120012033, JCYJ20240813112009013 and GJHZ20240218113603006); The Major Key Project of PCL for Experiments and Applications (PCL2024A08).
+Appendix A Literature Review
+We provide an overview of LLM-centric research of English Education presented in Figure 4.
+{forest}
+for tree= forked edges, grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=middle, font=, rectangle, line width=0.8pt, draw=output-black, rounded corners,align=left, minimum width=2em, s sep=6pt, l sep=8pt, , where level=1text width=0.2, where level=2text width=0.2font=, where level=3font=, where level=4font=, where level=5font=, [LLMs for English Education, middle,rotate=90,anchor=north,edge=output-black [LLM as Data Enhancer
+(Section 4),middle,anchor=west,edge=output-black, text width=0.20[Data Creation, middle, text width=0.14edge=output-black [ PFQS Li and Zhang (2024), MCQG-SRefine Yao et al. (2024),
+Lee et al. (2024c), BF-TC Liu et al. (2024d), MathVC Yue et al. (2024),
+EduAgent Xu et al.
+(2024), Generative Students Lu and Wang (2024) , leaf, text width=0.48edge=output-black] ] [Data Reformation, middle, text width=0.14edge=output-black [ Book2Dial Wang et al.
+(2024b), Slide2Lecture Zhang-Li et al.
+(2024),
+WikiDomains Asthana et al. (2024), Freyer et al. (2024),
+Day et al. (2025), Anthropological Prompting AlKhamissi et al. (2024),
+Liu et al. (2024a) , leaf, text width=0.48edge=output-black] ] [Data Annotation, middle, text width=0.14edge=output-black [ MEGAnno+ Kim et al. (2024), EDEN Siyan et al. (2024),
+PDFChatAnnotator Tang et al. (2024), Coannotating Li et al. (2023) , leaf, text width=0.48edge=output-black] ] ] [LLM as Task Predictor
+(Section 5),middle,anchor=west,edge=output-black, text width=0.20[Discriminative, middle, text width=0.14edge=output-black [ TOREE Zhuang et al. (2024), LLM-KT Zhan et al. (2024),
+CLST Jung et al. (2024), Diallogue-KT Scarlatos and Lan (2024),
+Neshaei et al. (2024), Mizumoto and Eguchi (2023), Song et al. (2025),
+Seßler et al. (2024), Syamkumar et al. (2024) , leaf, text width=0.48edge=output-black] ] [Generative, middle, text width=0.14edge=output-black [ RECIPE Han et al. (2023a), EXPECT Fei et al. (2023),
+GEE Song et al. (2024b), EXCGEC Ye et al. (2024),
+FELT Borges et al. (2024), LLM-as-a-tutor Han et al. (2024),
+SocraticLM Liu et al. (2024b), Stamper et al. (2024),
+Favero et al.
+(2024), Xiao et al.
+(2024), Scarlatos et al.
+(2025) , leaf, text width=0.48edge=output-black] ] [Mixed, middle, text width=0.14edge=output-black [ FABRIC Han et al.
+(2023b), eRevise+RF Liu et al.
+(2025),
+ReaLMistake Kamoi et al. (2024), Stahl et al. (2024b), Lu et al. (2024) , leaf, text width=0.48edge=output-black] ] ] [LLM-empowered Agent
+(Section 6),middle,anchor=west,edge=output-black, text width=0.20[Abilities, middle, text width=0.14edge=output-black [ FOKE Hu and Wang (2024), KnowEdit Zhang et al. (2024a),
+PedCoT Jiang et al. (2024b), LessonPlanner Fan et al. (2024),
+MemoryBank Zhong et al. (2024), vNMF Jiang et al. (2024a),
+ChatTutor Chen et al. (2024), LHP Sonkar et al. (2024),
+Razafinirina et al. (2024), Zhang et al. (2024a),
+Hu et al.
+(2024), Zheng et al.
+(2024), Shi et al.
+(2025) , leaf, text width=0.48edge=output-black] ] [Applications, middle, text width=0.14edge=output-black [ SimClass Zhang et al.
+(2024b), MathVC Yue et al.
+(2024),
+BIPED Kwon et al. (2024), LLM-as-a-tutor Han et al. (2024),
+KORLINGS Lee et al. (2024a), Liu et al. (2024d) , leaf, text width=0.48edge=output-black] ] ] ]
+Figure 4: An overview of LLM-centric research of FLE.
+Appendix B Future Directions
+Establishing Robust Evaluation Frameworks.
+A significant challenge in leveraging LLMs for English Education is the current lack of widely accepted and easily implementable evaluation frameworks to assess the quality of LLM-based teaching interactions and outcomes.
+Existing metrics often focus on linguistic correctness or task completion Tan et al.
+(2024); Macina et al.
+(2025) rather than pedagogical efficacy or impact on learning Chiang et al.
+(2024).
+Future work should prioritize the development of standardized evaluation methodologies, including comprehensive benchmarks and nuanced metrics that capture both the accuracy of linguistic information and the pedagogical value of LLM interventions.
+This will be essential for comparing different systems and guiding iterative improvements.
+Integrating with Modern Standardized Educational Frameworks.
+English language learning is often governed by established standards and frameworks, such as the Common European Framework of Reference for Languages (CEFR) [^1] https://www.coe.int/en/web/portal/home or Common Core State Standards (CCSS) [^2] https://corestandards.org/.
+For LLM-based tools to be truly effective and gain acceptance, their outputs and interaction patterns should align with these existing frameworks.
+Future technical development should focus on enabling LLMs to reference, interpret, and operate consistently within these standards Nicholls et al.
+(2024); Imperial et al.
+(2024).
+This includes generating proficiency-level-appropriate content, providing feedback that corresponds to specific framework descriptors, and assisting learners in achieving standardized learning objectives, thereby enhancing usability, conformity, and trustworthiness among educators and learners.
+Fostering Human-AI Collaboration in Pedagogy.
+While LLMs offer transformative potential, it is unlikely they will completely replace human teachers in English Education in the foreseeable future.
+Instead, the most promising path involves developing sophisticated human-AI collaborative educational technologies KIM et al.
+(2025).
+Future research should explore how LLMs can best function as assistive tools that augment, rather than supplant, the capabilities of human educators Shojaei et al.
+(2025).
+This includes designing intuitive interfaces for teachers to guide, customize, and oversee LLM-driven activities, investigating teachers’ perspectives on integrating LLMs into their practice, and defining technical benchmarks for when an LLM possesses sufficient acquired skills to reliably assist teachers.
+The focus must be on a synergistic model where LLMs handle scalable tasks while human teachers provide the crucial elements of empathy, nuanced understanding, and holistic student development.
+Appendix C Alternative Views
+While this paper supports the use of LLMs in English Education, it is essential to consider alternative perspectives. Below, we discuss two key opposing views and provide counterarguments.
+C.1 Task-Specific or Language-Specific Models as Better Alternatives
+Some argue that specialized or language-specific models, including classical ML systems with carefully engineered features, can outperform general-purpose LLMs in narrowly defined tasks (e.g., phonetics or grammar drills Fang et al.
+(2023)).
+By focusing on limited objectives, such models avoid the computational overhead and potential inaccuracies of LLMs, which aim to handle a broader range of inputs and contexts Shen et al.
+(2024a).
 Counterargument.
-While specialized models may excel in isolated tasks, they lack the flexibility required for comprehensive FLE, which involves cultural nuances, conversations, and evolving learner needs.
-In contrast, LLMs can be fine-tuned for specific goals while still offering broader linguistic competence (Song et al., 2024a).
+While specialized models may excel in isolated tasks, they lack the flexibility required for comprehensive English Education, which involves cultural nuances, conversations, and evolving learner needs.
+In contrast, LLMs can be fine-tuned for specific goals while still offering broader linguistic competence Song et al.
+(2024a).
 Additionally, relying on multiple specialized models can be resource-intensive, whereas a well-configured LLM provides a unified framework that balances specialization and scalability.
-8.2 Concerns About Over-Reliance on LLMs
-Critics warn that over-reliance on LLMs may lead to problems such as generating misleading outputs (Nahar et al., 2024), reducing human interaction, and over-standardizing teaching methods.
+C.2 Concerns About Over-Reliance on LLMs
+Critics warn that over-reliance on LLMs may lead to problems such as generating misleading outputs Nahar et al.
+(2024), reducing human interaction, and over-standardizing teaching methods.
 These issues could undermine the interpersonal and motivational aspects of language learning.
@@
 LLMs can complement educators by automating repetitive tasks, allowing teachers to focus on individualized support and motivation.
-Advances in AI safety, such as feedback loops (Tong et al., 2024) and human-in-the-loop systems (Wu et al., 2022), can help minimize inaccuracies (Ho et al., 2024).
-Additionally, the fine-tuning capabilities of LLMs ensure adaptability, supporting diverse and inclusive learning experiences (Lee et al., 2024b).
-9 Conclusion
-This position paper emphasizes the transformative potential of LLMs in FLE, positioning them as valuable tutors to complement traditional teaching methods.
-Through their roles as data enhancers, task predictors, and agents, LLMs can provide adaptive learning experiences across the core skills of listening, speaking, reading, and writing.
-This paper encourages interdisciplinary exploration to FLE.
-Impact Statement
-This paper positions LLMs as promising tutors in FLE.
-By leveraging LLMs, we aim to enhance inclusivity, personalization, and overall accessibility.
-However, potential risks—such as bias, privacy issues, and over-reliance on AI tutoring—necessitate careful oversight, which have been discussed in the paper.
-In particular, ensuring equitable data representation, model usage transparency, and human agency preservation remains critical.
-While careful integration with human expertise is essential, the potential for LLMs to revolutionize FLE is significant, paving the way for more inclusive, efficient, and engaging educational practices.
-We encourage continuing dialogue and interdisciplinary collaboration to responsibly integrate LLMs into educational ecosystems and mitigate potential harms.
-Appendix A Literature Review
-We provide an overview of LLM-centric research of FLE presented in Figure 4.
-LLMs for Foreign Language Education
-LLM as Data Enhancer (Section 4) Data Creation PFQS (Li & Zhang, 2024), MCQG-SRefine (Yao et al., 2024), (Lee et al., 2024c), BF-TC (Liu et al., 2024d), MathVC (Yue et al., 2024), EduAgent (Xu et al., 2024), Generative Students (Lu & Wang, 2024) Data Reformation Book2Dial (Wang et al., 2024b), Slide2Lecture (Zhang-Li et al., 2024), WikiDomains (Asthana et al., 2024), (Freyer et al., 2024), (Day et al., 2025), Anthropological Prompting (AlKhamissi et al., 2024), (Liu et al., 2024a) Data Annotation EXCGEC (Ye et al., 2024), EDEN (Siyan et al., 2024), MEGAnno+ (Kim et al., 2024), PDFChatAnnotator (Tang et al., 2024), Coannotating (Li et al., 2023) LLM as Task Predictor (Section 5) Discriminative TOREE (Zhuang et al., 2024), LLM-KT (Zhan et al., 2024), CLST (Jung et al., 2024), Diallogue-KT (Scarlatos & Lan, 2024), (Neshaei et al., 2024), (Mizumoto & Eguchi, 2023), (Seßler et al., 2024), (Syamkumar et al., 2024) Generative EXPECT (Fei et al., 2023), GEE (Song et al., 2024b), EXCGEC (Ye et al., 2024) FELT (Borges et al., 2024), LLM-as-a-tutor (Han et al., 2024), (Stamper et al., 2024) SocraticLM (Liu et al., 2024b), (Favero et al., 2024), (Xiao et al., 2024) Mixed FABRIC (Han et al., 2023), (Stahl et al., 2024b), ReaLMistake (Kamoi et al., 2024), (Lu et al., 2024) LLM-empowered Agent (Section 6) Abilities FOKE (Hu & Wang, 2024), KnowEdit (Zhang et al., 2024a), (Zhang et al., 2024a), PedCoT (Jiang et al., 2024b), LHP (Sonkar et al., 2024), (Razafinirina et al., 2024), LessonPlanner (Fan et al., 2024), MemoryBank (Zhong et al., 2024), (Hu et al., 2024), vNMF (Jiang et al., 2024a), ChatTutor (Chen et al., 2024), (Zheng et al., 2024) Applications SimClass (Zhang et al., 2024b), MathVC (Yue et al., 2024), BIPED (Kwon et al., 2024), LLM-as-a-tutor (Han et al., 2024), KORLINGS (Lee et al., 2024a), (Liu et al., 2024d)
-Figure 4: An overview of LLM-centric research of FLE.
-Appendix B Four Phases of Research Roadmap
-Stage 1: Rule-based Models (1960s–1990s).
-Early solutions relied on handcrafted linguistic rules to process language in tightly constrained scenarios (Grosan et al., 2011; C Angelides & Garcia, 1993).
-Classical platforms like PLATO (Hart, 1981) and Systran (Toma, 1977) operated effectively for highly structured tasks (e.g., grammar drills) but struggled with complex, context-dependent interactions.
-Stage 2: Statistical Models (1990s–2010s).
-With the increased availability of digitized corpora, methods such as the early version of Google Translate (Och, 2006) and Dragon NaturallySpeaking (Blair, 1997) pioneered statistical pattern mining.
-These approaches leveraged large datasets to infer linguistic rules probabilistically, improving scalability yet still lacking deeper semantic understanding.
-Stage 3: Neural Models (2010s–2020s).
-The advent of deep learning architectures (e.g., RNNs (Yu et al., 2019) and Transformers (Vaswani, 2017)) enabled more robust context modeling, sparking transformative applications like Grammarly (Fitria, 2021) and Duolingo (Vesselinov & Grego, 2012).
-These systems offered enhanced personalization and feedback, significantly augmenting learners’ writing and reading comprehension.
-Stage 4: Large Language Models (2020s–Present).
-Today’s LLMs (e.g., ChatGPT (Achiam et al., 2023)) combine massive pre-training corpora with generative capabilities, achieving impressive results in multi-turn dialogue, individualized scaffolding, and multimodal integration.
-Tools such as Khanmigo (Anand, 2023) demonstrate LLMs’ potential for real-time conversational practice, dynamic content creation, and inclusive educational support at scale.
+Advances in AI safety, such as feedback loops Tong et al.
+(2024) and human-in-the-loop systems Wu et al.
+(2022), can help minimize inaccuracies Ho et al.
+(2024).
+Additionally, the fine-tuning capabilities of LLMs ensure adaptability, supporting diverse and inclusive learning experiences Lee et al.
+(2024b).